### v1

# Charged spherically symmetric and slowly rotating charged black hole solutions in bumblebee gravity

Abstract
In this paper, we present charged spherically symmetric black hole solutions and slowly rotating charged solutions in bumblebee gravity with and without a cosmological constant. The static spherically symmetric solutions describe the Reissner-Nordström-like black hole and Reissner-Nordström-(anti) de Sitter-like black hole, while the stationary and axially symmetric soltuions describe Kerr-Newman-like black hole and Kerr-Newman-(anti) de Sitter-like black hole. We utilize the Hamilton-Jacobi formalism to study the shadows of the black holes. Additionally, we investigate the effect of the electric charge and Lorentz-violating parameters on the radius of the shadow reference circle and the distortion parameter. We find that the radius of the reference circle decreases with the Lorentz-violating parameter and charge parameter, while the distortion parameter increases with the Lorentz-violating parameter and the charge parameter.
Black Holes, Lorentz Symmetry Breaking,Black Hole Shadow
I INTRODUCTION
In the area of classical physics, gravitational phenomena are well-explained by general relativity, which has endured rigorous experimental and theoretical validation. On the quantum front, the Standard Model of particle physics has been successfully described the other three fundamental interactions. These two theories offer a comprehensive and understanding of the natural world. Therefore, the quest to unify these two theories is fundamental, and its achievement will inevitably lead us to a deeper comprehension of nature.
Several quantum gravity theories have been proposed Birrell and Davies (1984); Maldacena (1998); Aharony et al. (2000); Gubser et al. (1998); Alfaro et al. (2000, 2002); Rovelli and Smolin (1990), but it is needed the experiments at the Planck scale ($\sim 10^{19}$ GeV) to observe the quantum effect, this is beyond the ability of current experiments. However, some quantum gravity theories assume that Lorentz symmetry might be broken in the gravitational UV regime which is at our presently accessible low-energy scales. Stemming from the concept of spontaneous Lorentz symmetry breaking (LSB) in string theory Kostelecky and Samuel (1989a); Kostelecky (2004), the Standard Model extension was introduced as a suitable framework for violating Lorentz symmetry Colladay and Kostelecky (1997, 1998); Kostelecky (2004). The Standard Model extension has been widely studied Colladay and Kostelecky (2001); Kostelecky and Lehnert (2001); Kostelecky and Mewes (2001); Colladay and McDonald (2009); Berger et al. (2016); Carroll et al. (1990); Andrianov and Soldati (1995, 1998); Lehnert and Potting (2004); Baeta Scarpelli (2012); Brito et al. (2013).
In 1989, Kostelecky and Samuel introduced the prototype of the bumblebee model Kostelecky and Samuel (1989b, a), a string-inspired framework featuring tensor-induced spontaneous Lorentz symmetry breaking. In the framework of bumblebee gravity, the spontaneous Lorentz violation arises from a potential $V(B^{\mu}B_{\mu})$ acting on an vector field $B_{\mu}$ Kostelecky (2004); Kostelecky and Tasson (2011). Not just in Minkowski’s spacetime, The bumblebee model can also be explored in various spacetime geometries, including Riemann, Riemann-Cartan spacetimes Bailey and Kostelecky (2006); Maluf et al. (2013); Li and Övgün (2020); Bluhm et al. (2008a); Escobar and Martín-Ruiz (2017); Maluf et al. (2015); Hernaski (2014); Carroll et al. (2009).
The pursuit of black hole solutions holds significance for any theory of gravity. In 2017, Casana et al. obtained an exact Schwarzschild-like black hole solution in the bumblebee gravity Casana et al. (2018). Subsequently, researchers have independently proposed exact solutions for traversable wormholes Övgün et al. (2019), Schwarzschild-anti-de Sitter (AdS)-like black holes Maluf and Neves (2021), slowly rotating Kerr-like black holes Ding et al. (2020), and other solutions Santos et al. (2015); Jha and Rahaman (2021); Filho et al. (2023); Xu et al. (2023); Ding et al. (2023) within bumblebee gravity. Moreover, in other scenarios within the framework of the Standard Model Extension, such as when considering the nonminimally coupled Kalb-Ramond field with a nonzero vacuum expectation value, certain black hole solutions have also been obtained Lessa et al. (2020); Kumar et al. (2020); Yang et al. (2023); Duan et al. (2023); Do and Kao (2020); Liu et al. (2024a). And the properties of the blakc holes were also studied Guo et al. (2023); Du et al. (2024); Kuang and Övgün (2022); Oliveira et al. (2019); Liu et al. (2023); Gomes et al. (2020); Kanzi and Sakallı (2019); Güllü and Övgün (2022); Oliveira et al. (2021); Kanzi and Sakallı (2021); Hosseinifar et al. (2024); Liu et al. (2024b, a).
In 1966, Synge initiated the examination of the shadow cast by a Schwarzschild black hole Synge (1966). Subsequently, Luminet Luminet (1979) studied the influence of a thin accretion disk on the black hole shadow, while Bardeen Bardeen (1973) investigated the shadow of a Kerr black hole. Over the decades, significant progress has been made in the study of black hole shadows de Vries (1999); Takahashi (2005); Bambi et al. (2012); Atamurotov et al. (2013); Perlick et al. (2018); Meng et al. (2023); Sui et al. (2023). Recently, direct imaging of the black holes M87* at the core of the Virgo A galaxy and Sgr A* at the core of the Milky Way galaxy has been achieved by the sub-millimeter Event Horizon Telescope (EHT) employing very-long baseline interferometry Akiyama et al. (2019a, b, c, d); Akiyama et al. (2022). Therefore, investigating the black hole shadow is highly desirable.
Our aim in this paper is to obtain Charged spherically symmetric and slowly rotating charged black hole solutions. Subsequently, using the equations governing the motion of photons, we analyze the shadows cast by rotating black holes, exploring the impact of Lorentz violation on the size and deformation of these shadows. Calculations are then conducted to determine the radius and distortion observables.
This paper is structured as follows: Sec. II provides a review of the Einstein-Bumblebee theory. In Sec. III, we present the charged bumblebee black hole solution in both the presence and absence of a cosmological constant. And in Sec. IV, we extend this solution to the case of rotating black holes. Sec. V is devoted to investigating the shadow of rotating black holes and computing relevant parameters. Finally, Sec. VI presents a summary and discussion of the work.
II EINSTEIN-BUMBLEBEE THEORY
As discussed in the preceding section, the bumblebee model extends general relativity by adding a vector field called bumblebee field which couples with gravity nonminimally. The bumblebee vector field $B_{\mu}$ obtains a nonzero vacuum expectation value through a specified potential, resulting in the spontaneous breaking of Lorentz symmetry in the gravitational sector. The action is described by  Kostelecky (2004)
$S$ $=$ $\int d^{4}x\sqrt{-g}\bigg{[}\frac{1}{2\kappa}\left(R-2\Lambda\right)+\frac{\xi}{2\kappa}B^{\mu}B^{\nu}R_{\mu\nu}-\frac{1}{4}B_{\mu\nu}B^{\mu\nu}-V(B^{\mu}B_{\mu}\pm b^{2})\bigg{]}$ (1)
$+\int d^{4}x\sqrt{-g}\mathcal{L}_{M},$
where $\Lambda$ denotes the cosmological constant, $\kappa=\frac{8\pi G}{c^{4}}$ is the gravitational coupling constant, and $\xi$ represents the nonminimal coupling constant between gravity and the bumblebee field. Analogous to the electromagnetic field, the bumblebee field strength is given by
$B_{\mu\nu}=\partial_{\mu}B_{\nu}-\partial_{\nu}B_{\mu}.$ (2)
The potential $V$ provides a nonvanishing vacuum expectation value for $B_{\mu}$ and could take the functional form of $V(B^{\mu}B_{\mu}\pm b^{2})$, implying that the condition $B^{\mu}B_{\mu}=\mp b^{2}$ must be satisfied. In essence, the field $B_{\mu}$ acquires a nonvanishing vacuum expectation value $\left\langle B_{\mu}\right\rangle=b_{\mu}$, where the vector $b_{\mu}$ is a function of the spacetime coordinates and satisfies $b^{\mu}b_{\mu}=\mp b^{2}=\mathrm{const}$. For convenience, we define $X=B^{\mu}B_{\mu}\pm b^{2}$ and $V^{\prime}=\frac{\partial V}{\partial X}$, which will be used in the subsequent discussions.
We consider the matter field to be an electromagnetic field nonminimally coupled with the bumblebee vector field Lehum et al. (2024). Its Lagrangian density expression reads as
$\mathcal{L}_{M}$ $=$ $\frac{1}{2\kappa}(F^{\mu\nu}F_{\mu\nu}+{\gamma}B^{\mu}B_{\mu}F^{\alpha\beta}F_{\alpha\beta}),$ (3)
where the electromagnetic tensor related to the electromagnetic field is
$F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu},$
$A=(\phi(r),0,0,0),$ (4)
and ${\gamma}$ presents the coupling coefficient. The gravitational field equations in the framework of bumblebee gravity can be derived by varying the action (1) with respect to the metric tensor $g^{\mu\nu}$:
$G_{\mu\nu}+\Lambda g_{\mu\nu}=\kappa T^{B}_{\mu\nu}+\kappa T^{M}_{\mu\nu},$ (5)
where
$T^{B}_{\mu\nu}$ $=$ $\frac{\xi}{\kappa}\left[\frac{1}{2}B^{\alpha}B^{\beta}R_{\alpha\beta}g_{\mu\nu}-B_{\mu}B^{\alpha}R_{\alpha\nu}-B_{\nu}B^{\alpha}R_{\alpha\mu}\right.+\frac{1}{2}\nabla_{\alpha}\nabla_{\mu}\left(B^{\alpha}B_{\nu}\right)$ (6)
$+\frac{1}{2}\nabla_{\alpha}\nabla_{\nu}\left(B^{\alpha}B_{\mu}\right)\left.-\frac{1}{2}\nabla^{2}\left(B_{\mu}B_{\nu}\right)-\frac{1}{2}g_{\mu\nu}\nabla_{\alpha}\nabla_{\beta}\left(B^{\alpha}B^{\beta}\right)\right]$
$+2V^{\prime}B_{\mu}B_{\nu}+B_{\mu}^{\ \alpha}B_{\nu\alpha}-\left(V+\frac{1}{4}B_{\alpha\beta}B^{\alpha\beta}\right)g_{\mu\nu},$
and
$T^{M}_{\mu\nu}=\frac{1}{2\kappa}\left[(1+\gamma b^{2})(2F_{\mu\alpha}F_{\nu}^{\alpha}-\frac{1}{2}g_{\mu\nu}F^{\alpha\beta}F_{\alpha\beta})+\gamma B_{\mu}B_{\nu}F^{\alpha\beta}F_{\alpha\beta}\right].$ (7)
For computational convenience, we can express the field equations in the following form:
$R_{\mu\nu}=\Lambda g_{\mu\nu}+\kappa\mathcal{T}_{\mu\nu},$ (8)
where
$\mathcal{T}^{M}_{\mu\nu}$ $=$ ${T}^{M}_{\mu\nu}-\frac{1}{2}{T}^{M}g_{\mu\nu}$
$\mathcal{T}^{B}_{\mu\nu}$ $=$ ${T}^{B}_{\mu\nu}-\frac{1}{2}{T}^{B}g_{\mu\nu}$ (9)
and
$\kappa\mathcal{T}_{\mu\nu}$ $=$ $\kappa\mathcal{T}^{M}_{\mu\nu}+\kappa\mathcal{T}^{B}_{\mu\nu}$ (10)
$=$ $\kappa\left(T^{M}_{\mu\nu}-\frac{1}{2}g_{\mu\nu}T^{M}\right)+\kappa\left[V^{\prime}\left(2B_{\mu}B_{\nu}-b^{2}g_{\mu\nu}\right)+B_{\mu}^{\ \alpha}B_{\nu\alpha}+Vg_{\mu\nu}-\frac{1}{4}B_{\alpha\beta}B^{\alpha\beta}g_{\mu\nu}\right]$
$+\xi\left[\frac{1}{2}B^{\alpha}B^{\beta}R_{\alpha\beta}g_{\mu\nu}-B_{\mu}B^{\alpha}R_{\alpha\nu}-B_{\nu}B^{\alpha}R_{\alpha\mu}+\frac{1}{2}\nabla_{\alpha}\nabla_{\mu}\left(B^{\alpha}B_{\nu}\right)\right.+\frac{1}{2}\nabla_{\alpha}\nabla_{\nu}\left(B^{\alpha}B_{\mu}\right)$
$\left.-\frac{1}{2}\nabla^{2}\left(B_{\mu}B_{\nu}\right)\right].$
Similarly, by varying the action (1) with respect to the bumblebee vector field and the electromagnetic field, we can obtain the equations of motion for the corresponding fields:
$\nabla_{\mu}B^{\mu\nu}-2\left(V^{\prime}B^{\nu}-\frac{\xi}{2\kappa}B_{\mu}R^{\mu\nu}-\frac{1}{2\kappa}\gamma B^{\nu}F^{\alpha\beta}F_{\alpha\beta}\right)$ $=$ $0,$ (11)
$\nabla_{\mu}\left(F^{\mu\nu}+\gamma B^{\alpha}B_{\alpha}F^{\mu\nu}\right)$ $=$ $0$ (12)
III CHARGED SPHERICALLY SYMMETRIC BLACK HOLE SOLUTIONS
We consider the metric ansatz for a static and spherically symmetric spacetime, which is expressed as
${d}{s}^{2}=-A(r){dt}^{2}+S(r){dr}^{2}+r^{2}{\leavevmode\nobreak\ d}\Omega^{2},$ (13)
where ${\leavevmode\nobreak\ d}\Omega^{2}={\leavevmode\nobreak\ d}\theta^{2}+\sin^{2}\theta{d}\varphi^{2}$.
In the previous study conducted by Casana et al., they formulated an exact black hole solution without a cosmological constant and the matter field $A_{\mu}$ Casana et al. (2018). In the scenario where the bumblebee field $B_{\mu}$ maintains its vacuum expectation value $b_{\mu}$, the bumblebee field is given by Bertolami and Paramos (2005)
$B_{\mu}=b_{\mu}.$ (14)
Similar to Ref. Casana et al. (2018), we consider a spacelike background $b_{\mu}$ with the form
$b_{\mu}=\left(0,b_{r}(r),0,0\right).$ (15)
Utilizing the aforementioned condition $b_{\mu}b^{\mu}=b^{2}=const$, we can derive
$b_{r}(r)=b\sqrt{S(r)}.$ (16)
For convenience, we set $\ell=\xi b^{2}$ as Lorentz-violating parameter. By combining the spherically symmetric metric, we can obtain the specific field equations:
$\frac{(2+\ell)A^{\prime}(r)^{2}}{8A(r)S(r)}-\frac{(2+\ell)A^{\prime}(r)S^{\prime}(r)}{8S(r)^{2}}-\frac{(1+\ell)A^{\prime}(r)}{rS(r)}+\frac{(2+\ell)A^{\prime\prime}(r)}{4S(r)}$ (17)
$-\frac{\ell A(r)S^{\prime}(r)}{2rS(r)^{2}}-(1+2b^{2}\gamma)A(r)\phi^{\prime}(r)^{2}+\Lambda A(r)+\kappa(V^{\prime}b^{2}+V)A(r)$ $=$ $0,$ (18)
$\frac{A^{\prime}(r)S^{\prime}(r)}{4A(r)S(r)}-\frac{A^{\prime\prime}(r)}{2A(r)}+\frac{A^{\prime}(r)^{2}}{4A(r)^{2}}+\frac{S^{\prime}(r)}{rS(r)}$
$+\frac{2(1+2b^{2}\gamma)\phi^{\prime}(r)^{2}S(r)}{(2+3\ell)}-\frac{2\Lambda S(r)}{(2+3\ell)}-2\kappa\frac{(V^{\prime}b^{2}+V)S(r)}{(2+3\ell)}$ $=$ $0,$ (19)
$1-\frac{1+\ell}{S(r)}+\frac{rS^{\prime}(r)}{2S(r)^{2}}+\frac{\ell r^{2}A^{\prime\prime}(r)}{4A(r)S(r)}-\frac{\ell r^{2}A^{\prime}(r)S^{\prime}(r)}{8A(r)S(r)^{2}}-\frac{\ell r^{2}A^{\prime}(r)^{2}}{8A(r)^{2}S(r)}$
$-\frac{(1+\ell)rA^{\prime}(r)}{2A(r)S(r)}-(1+2b^{2}\gamma)\phi^{\prime}(r)^{2}r^{2}-\Lambda r^{2}-\kappa(V^{\prime}b^{2}+V)r^{2}$ $=$ $0,$ (20)
$\frac{A^{\prime\prime}(r)}{2A(r)}-\frac{A^{\prime}(r)^{2}}{4A(r)^{2}}-\frac{A^{\prime}(r)S^{\prime}(r)}{4A(r)S(r)}-\frac{2\gamma\phi^{\prime}(r)^{2}}{\xi\left(1+\ell\right)A(r)}-\frac{S^{\prime}(r)}{rS(r)}+\frac{2\kappa}{\xi}S(r)V^{\prime}$ $=$ $0,$ (21)
$\frac{d}{dr}\left[r^{2}\frac{F_{tr}}{\sqrt{A(r)S(r)}}\right]$ $=$ $0$ (22)
Next, we will solve the above equations and give spherically symmetric black hole solutions under different scenarios.
III.1 Case A: $V(X)=\frac{\lambda}{2}X^{2}$ and $\Lambda=0$
In the absence of the cosmological constant, similar to the work of Casana et al., we impose the vacuum conditions $V=0$ and $V^{\prime}=0$ Casana et al. (2018). An illustrative example of the potential satisfying these conditions is readily presented by a smooth quadratic form:
$V(X)=\frac{\lambda}{2}X^{2},$ (23)
where $\lambda$ is a constant.
Using Eq. (52), we can derive $F_{tr}=\sqrt{A(r)S(r)}\phi^{\prime}(r)$. Substituting the expressions of $F_{tr}$ here and $V$ in Eq. (23) into above conditions into the field equations (18) - (22), and considering the value of the coupling parameter $\gamma$ to be $\frac{\xi}{2+\ell}$, the field equations are given by:
$\frac{(2+\ell)A^{\prime}(r)^{2}}{8A(r)S(r)}-\frac{(2+\ell)A^{\prime}(r)S^{\prime}(r)}{8S(r)^{2}}-\frac{(1+\ell)A^{\prime}(r)}{rS(r)}$
$+\frac{(2+\ell)A^{\prime\prime}(r)}{4S(r)}-\frac{\ell A(r)S^{\prime}(r)}{2rS(r)^{2}}-\frac{A(r)\phi^{\prime}(r)^{2}}{\left(1+\ell\right)}$ $=$ $0,$ (24)
$\frac{A^{\prime}(r)S^{\prime}(r)}{4A(r)S(r)}-\frac{A^{\prime\prime}(r)}{2A(r)}+\frac{A^{\prime}(r)^{2}}{4A(r)^{2}}+\frac{S^{\prime}(r)}{rS(r)}+\frac{2\phi^{\prime}(r)^{2}S(r)}{(2+\ell)}$ $=$ $0,$ (25)
$1-\frac{1+\ell}{S(r)}+\frac{rS^{\prime}(r)}{2S(r)^{2}}+\frac{\ell r^{2}A^{\prime\prime}(r)}{4A(r)S(r)}-\frac{\ell r^{2}A^{\prime}(r)S^{\prime}(r)}{8A(r)S(r)^{2}}-\frac{\ell r^{2}A^{\prime}(r)^{2}}{8A(r)^{2}S(r)}$
$-\frac{(1+\ell)rA^{\prime}(r)}{2A(r)S(r)}-\frac{(2+3\ell)\phi^{\prime}(r)^{2}r^{2}}{(2+\ell)}$ $=$ $0,$ (26)
$\frac{A^{\prime\prime}(r)}{2A(r)}-\frac{A^{\prime}(r)^{2}}{4A(r)^{2}}-\frac{A^{\prime}(r)S^{\prime}(r)}{4A(r)S(r)}-\frac{\phi^{\prime}(r)^{2}}{\left(1+\ell\right)A(r)}-\frac{S^{\prime}(r)}{rS(r)}$ $=$ $0,$ (27)
$\frac{d}{dr}[r^{2}\phi^{\prime}(r)]$ $=$ $0,$ (28)
by simplifying Eq. (24) and Eq. (25), we obtain:
$[S(r)A(r)]^{\prime}=0,$ (29)
which implies that $S(r)=\frac{C_{1}}{A(r)}$, where $C_{1}$ is a constant. Similar to the Schwarzschild-like black hole solution in bumblebee gravity, we set the value of $C_{1}$ to be $1+\ell$ Casana et al. (2018). Building upon this, we solve the modified Maxwell’s equations to obtain: $\phi(r)=\frac{C_{2}}{r}+C_{3}$. Here, we set $C_{2}=Q_{0}$ and $C_{3}=0$ (i.e., $\phi(r)=\frac{Q_{0}}{r}$), and we ultimately arrive at the final solution:
$A(r)$ $=$ $1-\frac{2M}{r}+\frac{2(1+\ell)Q_{0}^{2}}{(2+\ell)r^{2}},$ (30)
$S(r)$ $=$ $\frac{1+\ell}{A(r)},$ (31)
$\phi(r)$ $=$ $\frac{Q_{0}}{r}.$ (32)
Through the modified Maxwell’s equations, the conserved current is modified to be $J^{\nu}=\nabla_{\mu}\left(F^{\mu\nu}+\gamma B^{\alpha}B_{\alpha}F^{\mu\nu}\right)$. Then, we can obtain the integration constants, the Maxwell’s equations was also modified.
[…]
$=$ $\left(1+b^{2}\frac{\xi}{(l+2)}\right)Q_{0}$
$=$ $\frac{2(1+\ell)}{2+\ell}Q_{0}.$
Here $\Sigma$ represents a 3-dimensional spacelike region with the induced metric $\gamma_{ij}^{(3)}$, the boundary $\partial\Sigma$ is a two-sphere located at spatial infinity with the induced metric $\gamma_{ij}^{(2)}=r^{2}{\leavevmode\nobreak\ d}\Omega^{2}$, while $n_{\mu}=(1,0,0,0)$ and $\sigma_{\mu}=(0,1,0,0)$ denote the unit normal vectors associated with $\partial\Sigma$ and $\Sigma$, respectively. The solution of $A(r)$ can also be rewritten as
$A(r)=1-\frac{2M}{r}+\frac{(2+\ell)Q^{2}}{2(1+\ell)r^{2}},$ (34)
This solution is similar to the Reissner-Nordström (RN) solution. As $\ell$ tends to 0, it recovers to the RN solution. Additionally, it is evident that the metric functions approach to $A(r)\rightarrow 1$ and $S(r)\rightarrow 1+\ell$ at infinity, indicating that the spacetime is not asymptotically Minkowski. Similar to the Reissner-Nordström black hole, this solution exhibits two horizons:
$r_{\pm}=M\pm\sqrt{M^{2}-\frac{2(\ell+1)}{\ell+2}Q_{0}^{2}}.$ (35)
It is obvious that the mass and charge parameter of the black hole should satisfy
$\frac{Q_{0}^{2}}{M^{2}}\leq\frac{(2+\ell)}{2(1+\ell)}.$ (36)
Another aspect of this solution is the physical singularity which can be studied through the Kretschmann scalar $K$, derived from the Riemann tensor:
$K$ $=$ $R_{\alpha\beta\delta\gamma}R^{\alpha\beta\delta\gamma}$ (37)
$=$ $\frac{4\left(\ell^{2}r^{2}+4lMr+12M^{2}\right)}{(\ell+1)^{2}r^{6}}-\frac{16Q_{0}^{2}\left((\ell+2)r(\ell r+12M)-14(\ell+1)Q_{0}^{2}\right)}{(\ell+1)(\ell+2)^{2}r^{8}}.$ (38)
From this we can see that, the singularity only appears at $r=0$ (assuming $0<\left|\ell\right|\ll 1$). This is similar to the RN black hole.
III.2 Case B: $V(X)=\frac{\lambda}{2}X$ and $\Lambda\neq 0$
Next, we consider the scenario involving a nonzero cosmological constant. Our purpose is to derive an analytical black hole solution in this context. The geometry of the solution is similar to the RN-Ads black hole or the RN-ds black hole. Based on the groundwork by R.V. Maluf et al. Maluf and Neves (2021), conditions are different from those in the previous discussion. The straightforward option for the potential can be a linear function:
$V(X)=\frac{\lambda}{2}X,$ (39)
where $\lambda$ is a Lagrange multiplier Bluhm et al. (2008b). Note that the equation of motion for the Lagrange multiplier ensures the vacuum condition $X=0$, which also ensures that $V=0$ for any field $\lambda$ on-shell. However, unlike the vacuum condition in Case A, $V^{\prime}\neq 0$ is valid when the field $\lambda$ is nonzero. Thus, the potential $V$ contributes to the field equations. The constraint for the Lagrange multiplier $\lambda$ is providedthe field Eqs (10) and (11) .
Combining the spherically symmetric metric and the form of the potential $V$, we can derive the specific field equations:
$\frac{(2+\ell)A^{\prime}(r)^{2}}{8A(r)S(r)}-\frac{(2+\ell)A^{\prime}(r)S^{\prime}(r)}{8S(r)^{2}}-\frac{(1+\ell)A^{\prime}(r)}{rS(r)}$
$+\frac{(2+\ell)A^{\prime\prime}(r)}{4S(r)}-\frac{\ell A(r)S^{\prime}(r)}{2rS(r)^{2}}-\frac{A(r)\phi^{\prime}(r)^{2}}{\left(1+\ell\right)}+\frac{\Lambda A(r)(2+\ell)}{2\ell}$ $=$ $0,$ (40)
$\frac{A^{\prime}(r)S^{\prime}(r)}{4A(r)S(r)}-\frac{A^{\prime\prime}(r)}{2A(r)}+\frac{A^{\prime}(r)^{2}}{4A(r)^{2}}+\frac{S^{\prime}(r)}{rS(r)}+\frac{2\phi^{\prime}(r)^{2}S(r)}{(2+\ell)}-\frac{\Lambda S(r)}{1+\ell}$ $=$ $0,$ (41)
$1-\frac{1+\ell}{S(r)}+\frac{rS^{\prime}(r)}{2S(r)^{2}}+\frac{\ell r^{2}A^{\prime\prime}(r)}{4A(r)S(r)}-\frac{\ell r^{2}A^{\prime}(r)S^{\prime}(r)}{8A(r)S(r)^{2}}-\frac{\ell r^{2}A^{\prime}(r)^{2}}{8A(r)^{2}S(r)}$
$-\frac{(1+\ell)rA^{\prime}(r)}{2A(r)S(r)}-\frac{(2+3\ell)\phi^{\prime}(r)^{2}r^{2}}{(2+\ell)}-\frac{\Lambda(2+\ell)r^{2}}{2(1+\ell)}$ $=$ $0,$ (42)
$\frac{A^{\prime\prime}(r)}{2A(r)}-\frac{A^{\prime}(r)^{2}}{4A(r)^{2}}-\frac{A^{\prime}(r)S^{\prime}(r)}{4A(r)S(r)}-\frac{\phi^{\prime}(r)^{2}}{\left(1+\ell\right)A(r)}-\frac{S^{\prime}(r)}{rS(r)}+\frac{\Lambda S(r)}{1+\ell}$ $=$ $0,$ (43)
$\frac{d}{dr}[r^{2}\phi^{\prime}(r)]$ $=$ $0$ (44)
An analytical solution is be possible if and only if
$\Lambda=\frac{\kappa\lambda}{\xi}(1+\ell).$ (45)
Under the above condition, we can obtain
$A(r)$ $=$ $1-\frac{2M}{r}+\frac{2(1+\ell)Q_{0}^{2}}{(2+\ell)r^{2}}-\frac{1}{3}(1+\ell)r^{2}\Lambda_{e},$ (46)
$S(r)$ $=$ $\frac{1+\ell}{1-\frac{2M}{r}+\frac{2(1+\ell)Q_{0}^{2}}{(2+\ell)r^{2}}-\frac{1}{3}(1+\ell)r^{2}\Lambda_{e}},$ (47)
$\phi(r)$ $=$ $\frac{Q_{0}}{r}.$ (48)
For convenience, we set $\Lambda_{e}=\frac{\kappa\lambda}{\xi}$ as an effective cosmological constant, which satisfies the relation $R_{\mu\nu}=\Lambda_{e}g_{\mu\nu}$ at the boundary of the spacetime and is independent of the Lorentz-violating parameter.
As $\ell$ approaches to 0, this solution becomes RN-(A)ds black hole solution. Similarly, we compute the Kretschmann scalar in order to examine the singularity study the singularity. The expression for $K$ is
$K$ $=$ $\frac{8}{3}\Lambda_{e}^{2}+\frac{l}{(l+1)r^{2}}\Lambda_{e}-\frac{16Q_{e}^{2}\left((\ell+2)r(\ell r+12M)-14(\ell+1)Q_{e}^{2}\right)}{(\ell+1)(\ell+2)^{2}r^{8}}$ (49)
$+$ $\frac{4\left(\ell^{2}r^{2}+4\ell Mr+12M^{2}\right)}{(\ell+1)^{2}r^{6}},$
which shows that, the RN-(A)ds-like singularity is only present at $r=0$ (assuming $0<\left|\ell\right|\ll 1$).
IV SlOWLY ROTATING CHARGED BLACK HOLE SOLUTIONS
In this section, we will give slowly rotaing charged solutions through solving Einstein-bumblebee equations in both the cases with and without a cosmological constant. The forms of these solutions are very similar to the Kerr-Newman (KN) and Kerr-Newman (KN)-(A)ds solutions.
In 1963, Kerr discovered the rotating black hole solution Kerr (1963). A few years later, Newman obtained rotating charged black hole solution Newman et al. (1965). In bumblebee gravity, some rotating solutions were also provided Ding et al. (2020); Ding and Chen (2021); Jha and Rahaman (2021). Here, we consider slowly rotating charged black hole solution with and without a cosmological constant. The general form of the corresponding metric is given by
$ds^{2}=-A(r,\theta)dt^{2}+S(r,\theta)dr^{2}+2F(r)H(\theta)adtd\phi+\rho(r,\theta)^{2}d\theta^{2}+h(r,\theta)^{2}\sin^{2}\theta d\phi^{2}.$ (50)
The bumblee field and the $U(1)$ vector field are assumed as
$b_{\mu}$ $=$ $(0,b_{r}(r,\theta),0,0),$ (51)
$F_{\mu\nu}$ $=$ $\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu},$
$A_{\mu}$ $=$ $(A_{0}(r,\theta),0,0,A_{\phi}(r,\theta)),$ (52)
IV.1 Case A: $V(X)=\frac{\lambda}{2}X^{2}$ and $\Lambda=0$
It is noted that when the Lorentz-violating parameter is 0, this solution should reduce to the Kerr-Newman solution. Similarly, when $a=0$, we should obtain the spherically symmetric black hole solution derived in the previous section. Therefore, we consider the product of $a$ and $\ell$ is a small quantity.
Ding et al. obtained a slowly rotating black hole solution within bumblebee gravity  Ding et al. (2020); Ding and Chen (2021). Following their approach, we obtained a slowly rotating charged black hole solution within bumblebee gravity. The metric in Boyer-Lindquist coordinates reads:
$ds^{2}$ $=$ $-\frac{\Delta_{r}}{\rho^{2}}\Big{(}\mathrm{d}t-a\sqrt{1+\ell}\sin^{2}\theta\mathrm{d}\phi\Big{)}^{2}+(1+\ell)\frac{\rho^{2}}{\Delta_{r}}\mathrm{d}r^{2}+{\rho^{2}}\mathrm{d}\theta^{2}$ (53)
$\quad+\frac{\sin^{2}\theta}{\rho^{2}}\Big{(}a\sqrt{1+\ell}\mathrm{d}t-(r^{2}+a^{2}(1+\ell))\mathrm{d}\phi\Big{)}^{2},$
$A_{0}(r,\theta)$ $=$ $-\frac{Q_{0}r}{\rho^{2}},$ (54)
$A_{\phi}(r,\theta)$ $=$ $\frac{Q_{0}a\sqrt{1+\ell}r\sin^{2}\theta}{\rho^{2}},$ (55)
$b_{r}(r)$ $=$ $b\rho\sqrt{\frac{\left(1+\ell\right)}{\Delta_{r}}},$ (56)
where
$\Delta_{r}$ $=$ $(1+\ell)a^{2}+r^{2}-2Mr+\frac{2(1+\ell)Q_{0}^{2}}{(2+\ell)},$
$\rho$ $=$ $\sqrt{r^{2}+(1+\ell)a^{2}\text{cos}^{2}\theta}.$ (57)
When $\ell\rightarrow 0$, it recovers to the usual Kerr-Newman metric and when $a\rightarrow 0$, it becomes to metric (13) with Eqs. (30) and (31). By solving $\Delta_{r}=0$, we can obatin the radius of the black hole’s event horizon
$r_{\pm}=M\pm\sqrt{M^{2}-\frac{2(\ell+1)Q_{0}^{2}+(\ell+2)a^{2}}{\ell+2}}.$ (58)
It becomes evident that the mass and spin parameters of the black hole must satisfy.
$\frac{2(\ell+1)Q_{0}^{2}+(\ell+2)a^{2}}{\ell+2}\leq M^{2}.$ (59)
Next, we discuss the small perturbations introduced by this rotating solution to the field equations, and impose certain constraints on the parameters through these perturbations. We set
$\Delta_{\mu\nu}=$ $R_{\mu\nu}$ $-\kappa\left[V^{\prime}\left(2B_{\mu}B_{\nu}+b^{2}g_{\mu\nu}\right)+B_{\mu}^{\ \alpha}B_{\nu\alpha}+Vg_{\mu\nu}\frac{1}{4}B_{\alpha\beta}B^{\alpha\beta}g_{\mu\nu}\right]$ (60)
-- $\kappa(T^{M}_{\mu\nu}-\frac{1}{2}g_{\mu\nu}T^{M})-\xi\left[\frac{1}{2}B^{\alpha}B^{\beta}R_{\alpha\beta}g_{\mu\nu}-B_{\mu}B^{\alpha}R_{\alpha\nu}-B_{\nu}B^{\alpha}R_{\alpha\mu}\right.$
$+$ $\frac{1}{2}\nabla_{\alpha}\nabla_{\mu}\left(B^{\alpha}B_{\nu}\right)+\frac{1}{2}\nabla_{\alpha}\nabla_{\nu}\left(B^{\alpha}B_{\mu}\right)\left.-\frac{1}{2}\nabla^{2}\left(B_{\mu}B_{\nu}\right)\right],$
naturally,for finite $Q_{0}$, when the field equations hold, we have $\Delta_{\mu\nu}=0$. Substituting the metric into the right hand side of Eq. (60), our calculations yield:
$\Delta_{\mu\nu}=\left\{\begin{aligned} \mathcal{O}(a^{2}\ell)&,&\quad\mu\nu=tt,rr,\phi\phi,\theta\theta,r\theta,t\phi\\ 0&,&\quad\text{otherwise},\end{aligned}\right.$ (61)
and the remaining components of $\bar{R}_{\mu\nu}=0$ are all zero. At the same time, we derive the equations of motion for the bumblebee field and the electromagnetic field. Only the $r$ component nonvanishing:
$\left(\nabla_{\mu}B^{\mu r}-2(V^{\prime}B^{r}-\frac{\xi}{2\kappa}B_{\mu}R^{\mu r}-\gamma B^{r}F^{\alpha\beta}F_{\alpha\beta})\right)b$ $=$ $\mathcal{O}(a^{2}\ell),$
$\nabla_{\mu}\left(F^{\mu r}+\gamma B^{\alpha}B_{\alpha}F^{\mu r}\right)$ $=$ $\mathcal{O}(a^{2}\ell).$ (62)
Combining the above equations, we find that the rotating solution needs to satisfy the condition $\mathcal{O}(a^{2}\ell)$ being a small quantity. This implies that when considering this solution,$a$ and $\ell$ cannot both take large values simultaneously. Thus, for our solution, we need the condition that $a^{2}\ell$ is small quantities ($\frac{a^{2}\ell}{M^{2}}\ll 1$) for finite $Q_{0}$ , respectively.
IV.2 Case B: $V(X)=\frac{\lambda}{2}X$ and $\Lambda\neq 0$
Next, we consider the case of $V(X)=\frac{\lambda}{2}X$ and $\Lambda\neq 0$. Similar to the case A, we can get the following solution:
$ds^{2}$ $=$ $-\frac{\Delta_{r}}{\rho^{2}}\Big{(}\mathrm{d}t-\frac{a\sqrt{1+\ell}\sin^{2}\theta}{\Xi}\mathrm{d}\phi\Big{)}^{2}+(1+\ell)\frac{\rho^{2}}{\Delta_{r}}\mathrm{d}r^{2}+\frac{\rho^{2}}{\Delta_{\theta}}\mathrm{d}\theta^{2}$ (63)
$\quad+\frac{\sin^{2}\theta\Delta_{\theta}}{\rho^{2}}\Big{(}a\sqrt{1+\ell}\mathrm{d}t-\frac{r^{2}+a^{2}(1+\ell)}{\Xi}\mathrm{d}\phi\Big{)}^{2},$
$A_{0}(r,\theta)$ $=$ $-\frac{Q_{0}r}{\rho^{2}},$ (64)
$A_{\phi}(r,\theta)$ $=$ $\frac{Q_{0}ar\sin^{2}\theta}{\Xi\rho^{2}},$ (65)
$b_{r}(r)$ $=$ $b\rho\sqrt{\frac{\left(1+\ell\right)}{\Delta_{r}}},$ (66)
where
$\Delta_{\theta}$ $=$ $1+\frac{1}{3}(1+\ell)^{2}\Lambda_{e}a^{2}\text{cos}^{2}\theta,$
$\Delta_{r}$ $=$ $\left((1+\ell)a^{2}+r^{2}\right)\left(1-\frac{1}{3}(1+\ell)\Lambda_{e}r^{2}\right)-2Mr+\frac{2(1+\ell)Q_{0}}{(2+\ell)},$
$\Xi$ $=$ $1+\frac{1}{3}(1+\ell)^{2}\Lambda_{e}a^{2},$
$\rho$ $=$ $\sqrt{r^{2}+(1+\ell)a^{2}\text{cos}^{2}\theta}.$ (67)
From the form of the solution, it is a KN-(A)dS-like solution. Note that when the charge parameter $Q_{\text{0}}=0$, the solution is reduced to a slowly rotating Kerr-(A)dS-like black hole solution, which is not found in previous work. Otherwise, when $\ell=0$ and $a=0$, we get the usual KN-(A)dS black hole solution and the metric (13) with (46) and (47), respectively.
By calculating the errors in the field equations, we find that the results are the same as case A. Similarly, the rotating solution must satisfy the condition that $a^{2}\ell$ is a small quantity.
V SHADOW OBSERVABLES
The black hole shadow is formed by the strong gravitational field of the black hole. The shadow appears as a two-dimensional dark region to a distant observer. Shadows serve as one of the most intuitive pieces of evidence for the presence of black holes. Particularly, the unveiling of an image provides even stronger evidence for black hole’s existence Akiyama et al. (2019a, b, c, d); Akiyama et al. (2022). The shadow of the Kerr-like black hole in bumblebee gravity has been extensively studied in previous works Wang and Wei (2022); Ding et al. (2020); Kuang and Övgün (2022); Chen et al. (2020). In this section, we aim to investigate the shadow of the KR-AdS like black holes in the framework of Einstein-bumblebee gravity.
V.1 The photon orbit
In this section, we will derive the equations of motion of photons orbiting the black holes. From Eqs. (61) and (62), it is apparent that the field equations hold when the rotating black hole solution satisfies that $a\ell^{2}$ is a small quantity. Therefore, in the subsequent part, all the analysies are under this condition. The motion of a photon in the background of the black hole is given by the geodesics. However, the process of solving geodesic equations is very difficult. we choose to use the Hamilton-Jacobi approach to solve it. The Lagrangian for the photon is
$\mathcal{L}=\frac{1}{2}g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu},$ (68)
where $\dot{x}$ is defined as $\dot{x}^{\mu}=dx^{\mu}/d\lambda=u^{\mu}$, with the four-velocity of the photon $u^{\mu}$ and the affine parameter $\lambda$. To facilitate the study of photon orbits and the black hole shadow observed by a distant observer, we introduce two parameters:
$\xi=\frac{L_{z}}{E}\,,\quad\eta=\frac{K}{E^{2}}\,,$ (69)
where $E$, $L_{z}$, and $K$ represent the energy, the axial component of the angular momentum, and the Carter constant Carter (1968), respectively. The expressions of the energy and the $z$-component of the angular momentum is given by
$E=-\frac{\partial{\mathcal{L}}}{\partial\dot{t}}=-g_{\varphi t}\dot{\varphi}-g_{tt}\dot{t},$ (70)
$L_{z}=\frac{\partial{\mathcal{L}}}{\partial\dot{\varphi}}=g_{\varphi\varphi}\dot{\varphi}+g_{\varphi t}\dot{t}.$ (71)
Using the Hamilton-Jacobi equation, we can get the null geodesic equation of the photon around the slowly rotating black hole:
$\dot{t}$ $=$ $\frac{E}{\rho^{2}}\Bigg{[}\frac{\left(r^{2}+(1+\ell)a^{2}\right)\left[\left(r^{2}+(1+\ell)a^{2}\right)-\sqrt{1+\ell}a\xi\Xi\right]}{\Delta_{r}}$
$+\frac{\sqrt{1+\ell}a\left(\xi\Xi-\sqrt{1+\ell}a\sin^{2}\theta\right)}{\Delta_{\theta}}\Bigg{]},$
$\dot{\varphi}$ $=$ $\frac{E\Xi}{\rho^{2}}\left[\frac{\sqrt{1+\ell}a(\left(r^{2}+(1+\ell)a^{2}\right)-\sqrt{1+\ell}a\xi\Xi)}{\Delta_{r}}+\frac{\xi\Xi-\sqrt{1+\ell}a\sin^{2}\theta}{\sin^{2}\theta\Delta_{\theta}}\right],$ (73)
$(1+\ell)(\rho^{2}\dot{r})^{2}$ $=$ $E^{2}\left[\left[\left(r^{2}+(1+\ell)a^{2}\right)-\sqrt{1+\ell}a\xi\Xi\right]^{2}-\Delta_{r}\Big{[}\frac{\left(\sqrt{1+\ell}a-\xi\Xi\right)^{2}}{\Delta_{\theta}}+\eta\Big{]}\right]$ (74)
$\equiv$ $(1+\ell)\mathcal{R}(r),$
$\rho^{2}\dot{\theta}$ $=$ $E^{2}\left[\eta\Delta_{\theta}-\cos^{2}\theta\left(\frac{\xi^{2}\Xi^{2}}{\sin^{2}\theta}-(1+\ell)a^{2}\right)\right]\equiv\Theta(\theta).$ (75)
Figure 1: The effective potential $V_{\text{eff}}$ for the photon. The paramengts are set as $M=1,a=0.1$ , $Q_{0}=0.1$ , $l=0,0.1$ and $\xi_{c}$ is an the value of $\xi$ for unstable circular null orbit.
Photons with sufficiently large orbital angular momentum can escape from the black hole, while those with small angular momentum will fall into the black hole. Now we focus on the critical case, the circular null orbit, where the photons will neither escape from nor fall into the black hole. The radial motion can also be expressed as
$(\rho^{2}\dot{r})^{2}+V_{eff}=0,$ (76)
where the effective potential $V_{\text{eff}}$ reads
$V_{eff}=-\frac{E^{2}}{(1+\ell)}\left[\left[\left(r^{2}+(1+\ell)a^{2}\right)-\sqrt{1+\ell}a\xi\Xi\right]^{2}-\Delta_{r}\big{[}\frac{\left(\sqrt{1+\ell}a-\xi\Xi\right)^{2}}{\Delta_{\theta}}+\eta\big{]}\right].$ (77)
In Fig. 1, we depict the effective potential $V_{\text{eff}}$ for a rotating charged black hole with $\ell=0$ and $\ell=0.1$. From Fig. 1, we can observe that Lorentz-violating parameter greater than 0 can increase the effective potential Wei et al. (2015). For the unstable circular null orbit, the effective potential satisfies
${V}(r)=\frac{d{V}(r)}{dr}=0,\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \frac{d^{2}V(r)}{dr^{2}}<0,$ (78)
Then we can obtain
$\eta$ $=$ $\frac{r^{2}\left(16(1+\ell)a^{2}\Delta_{r}\Delta_{\theta}-\left(r\Delta_{r}^{\prime}-4\Delta_{r}\right)^{2}\right)}{(1+\ell)a^{2}\Delta_{\theta}\Delta_{r}^{\prime 2}},$ (79)
$\xi$ $=$ $\frac{\left(r^{2}+(1+\ell)a^{2}\right)\Delta_{r}^{\prime}-4r\Delta_{r}}{\sqrt{1+\ell}a\Xi\Delta_{r}^{\prime}}.$ (80)
(a)  $M=1,a=0.1,\ell=0.01,Q_{0}=0.2,\theta=\frac{\pi}{2}.$
(b)  $M=1,a=0.9,\ell=0.01,\Lambda_{e}=-0.1,Q_{0}=0.3.$
(c)  $M=1,a=0.1,\Lambda_{e}=-0.1,Q_{0}=0.2.$
(d)  $M=1,a=0.647,\ell=0.01,\Lambda_{e}=-0.2,\theta=\frac{\pi}{2}.$
Figure 2: The shadows for the KN-(A)dS-like black hole with different parameters.
V.2 Observables
We now aim to determine the apparent shape of the slowly rotating charged black hole’s shadow. The apparent shape of the black hole shadow for an observer far away from the black hole can be described using the celestial coordinates $x$ and $y$:
$x$ $=$ $\lim_{r\to\infty}\frac{rp^{(\theta)}}{p^{(t)}}=\sqrt{\eta+a^{2}\cos^{2}\theta-\xi^{2}\cot^{2}\theta},$ (81)
$y$ $=$ $\lim_{r\to\infty}\frac{-rp^{(\varphi)}}{p^{(t)}}=-\xi\csc\theta.$ (82)
The coordinate $x$ denotes the apparent perpendicular distance from the shape’s projection on the equatorial plane, while $y$ represents the apparent perpendicular distance of the shape from the axis of symmetry. The boundary curve of the shadow is determined by the unstable circular null orbit. Substituting the metric into Eqs. (79)and (80), we have
$\eta$ $=$ $\frac{4r^{2}\left(\left(a^{2}(\ell+1)+r^{2}\right)\left(1-\frac{1}{3}(\ell+1)r^{2}\Lambda_{e}\right)-2Mr+\frac{(1+\ell)Q_{0}^{2}}{(1+\frac{\ell}{2})}\right)}{\left(\frac{1}{3}(\ell+1)r\Lambda_{e}\left(a^{2}(\ell+1)+2r^{2}\right)+M-r\right)^{2}}$ (83)
$-\frac{r^{2}\left(-\frac{1}{3}a^{2}(\ell+1)r^{2}\Lambda_{e}+2a^{2}-\frac{3Mr}{1+\ell}+\frac{4Q_{0}^{2}}{(2+\ell)}+r^{2}\right)^{2}}{a^{2}\left(\frac{1}{3}a^{2}(\ell+1)^{2}\Lambda_{e}\cos^{2}(\theta)+1\right)\left(\frac{1}{3}(\ell+1)r\Lambda_{e}\left(a^{2}(\ell+1)+2r^{2}\right)+M-r\right){}^{2}},$
$\xi$ $=$ $\frac{\frac{1}{3}a^{2}(\ell+1)^{2}r\Lambda_{e}\left(a^{2}(\ell+1)+r^{2}\right)+a^{2}(\ell+1)(M+r)+r\left(r(r-3M)+\frac{4Q_{0}^{2}}{(2+\ell)}\right)}{a\sqrt{\ell+1}\left(\frac{1}{3}a^{2}(\ell+1)^{2}\Lambda_{e}+1\right)\left(\frac{1}{3}(\ell+1)r\Lambda_{e}\left(a^{2}(\ell+1)+2r^{2}\right)+M-r\right)}.$ (84)
(a)  $\theta=\frac{\pi}{2}$
(b)  $\theta=\frac{\pi}{4}$
(c)  $\theta=\frac{\pi}{6}$
Figure 3: The radius $R_{s}$ of the shadow against the Lorentz-violating parameter $\ell$.
(a)  $\theta=\frac{\pi}{2}$
(b)  $\theta=\frac{\pi}{4}$
(c)  $\theta=\frac{\pi}{6}$
Figure 4: The distortion parameter $\delta_{s}$ against the Lorentz-violating parameter $\ell$.
It is easy to see that both the coordinates $x$ and $y$ are the functions of $(M,Q_{eff},a,\Lambda_{e},\ell,\theta)$. We set $M=1$, and show the shapes of the shadow in Fig. 2. Similar to the Kerr-Newman-AdS black hole, the shadow radius increases with the negative effective cosmological costant $\Lambda_{e}$, as shown in Fig. 2(a). From Fig. 2(b) we can see that, the position of the shadow depends on the observation angle $\theta$. The Lorentz-violating parameter $\ell$ also affects the shaow size which can be seen from Fig. 2(c). Besides, the deformation of the shadow becomes more pronounced when the charge parameter $Q_{0}$ increases, which can be seen from Fig. 2(d).
(a)   $\theta=\frac{\pi}{2}$
(b)   $\theta=\frac{\pi}{4}$
(c)  $\theta=\frac{\pi}{6}$
Figure 5: The radius $R_{s}$ of the shadow against the charge parameter $Q_{0}$.
(a)  $\theta=\frac{\pi}{2}$
(b)  $\theta=\frac{\pi}{4}$
(c)  $\theta=\frac{\pi}{6}$
Figure 6: The distortion parameter $\delta_{s}$ against the charge parameter $Q_{0}$.
To effectively analyze astronomical observational data, it is essential to establish easily measurable and reliable astronomical observables. In this study, we will examine the following observables, the radius $R_{s}$ and the distortion parameter $\delta_{s}$ which are defined by Hioki and Maeda Hioki and Maeda (2009). The parameter $R_{s}$ represents the radius of the reference circle, which is expected to pass through the top point $(x_{t},y_{t})$, the bottom point $(x_{b},y_{b})$, and the right point $(x_{r},0)$ of the shadow. Meanwhile, the distortion parameter quantifies the distortion of the black hole shadow in comparison with the reference circle, and they can be expressed as
$R_{s}=\frac{(x_{t}-x_{r})^{2}+y_{t}^{2}}{2(x_{r}-x_{t})},\quad\delta_{s}=\frac{x_{l}-\tilde{x}_{l}}{R_{s}},$ (85)
where $(x_{l},0)$ and $(\tilde{x}_{l},0)$ represent the left points of the shadow and the reference circle, respectively.
For a given spin parameter, we illustrate the radius $R_{s}$ and the distortion parameter $\delta_{s}$ as functions of the Lorentz-violating parameter $\ell$ for $\theta=\frac{\pi}{2},\theta=\frac{\pi}{4},\theta=\frac{\pi}{6}$. The behavior of $R_{s}$ and $\delta_{s}$ is exhibited in Fig. 4 and Fig. 4. For fixed $a$ and $\theta$, the radius $R_{s}$ decreases with $\ell$, but $\delta_{s}$ increases with $\ell$ increases. Particularly noteworthy is the observation that for observers situated far from the equatorial plane, the shadow of a rapidly rotating black hole appears smaller compared to that of a slower one with the same $\ell$. However, this discrepancy diminishes as the observation angle increases, eventually resulting in a convergence of all curves as $\theta$ approaches $\frac{\pi}{2}$. Furthermore, as the observer progressively approaches the equatorial plane, the distortion of the shadow becomes increasingly pronounced.
Similarly, we investigate the relationship between the value of the charge parameter and $R_{s}$ and $\delta_{s}$ which is shown in Fig. 6 and Fig. 6. We find that $R_{s}$ decreases with the charge parameter, whereas $\delta_{s}$ increases with the charge parameter. Another notable observation in our study is that for small values of the spin parameter $a$, the deformation of the shadow is nearly imperceptible.
VI CONCLUSIONS
In this paper, by considering the coupling between the electromagnetic field and the bumblebee vector field, we obtained the RN-like black hole solutions and the RN-(A)dS-like black hole solutions in bumblebee gravity. When the Lorentz-violating parameter vanishes $(\ell=0)$, these solutions reduce to the RN and RN-(A)dS black hole solutions. In the stationary and axially symmetric scenario, we obtained solutions for slowly rotating charged black holes, i.e., the KN-(A)dS-like black hole, within the framework of bumblebee gravity. We further analyzed the series expansion of small quantities.
Because the nonminimal interaction term between the electromagnetic field and bumblebee field was introduced, the Maxwell’s equations was also modified. The charge term of the black hole solution is related to the Lorentz-violating parameter. When the electromagnetic field does not exist, our solutions are then reduced to another bumblebee black hole solutions in previous works.
Moreover, in the background of the KN-(A)dS-like black hole, we derived null geodesics using the Hamilton-Jacobi equation. Utilizing the conditions for unstable circular null orbits, we derived two conserved parameters, $\xi$ and $\eta$, which govern the photon’s motion. Using these parameters, we determined the celestial coordinates of the shadow. Subsequently, we investigated the size and deformation of the shadow. By varying the observation angle $\theta=\frac{\pi}{2},$ $\theta=\frac{\pi}{4},$ $\theta=\frac{\pi}{6}$, we calculated the shadow’s radius $R_{s}$ for different parameters. We observed that for fixed spin $a$ and Lorentz-violating parameter $\ell$, observers situated at lower latitudes consistently observe a larger shadow. Meanwhile, we found that the radius of the reference circle decreases with the Lorentz-violating parameter and electric charge , while the distortion parameter increases with the Lorentz-violating parameter and electric charge.
Acknowledgements
We are grateful to Tao-tao Sui and Ke Yang for useful discussions. This work was supported by the National Key Research and Development Program of China (Grant No. 2021YFC2203003), the National Natural Science Foundation of China (Grants No. 12205129, No. 12147166, and No. 12247101), the China Postdoctoral Science Foundation (Grant No. 2021M701529), the 111 Project (Grant No. B20063), the Major Science and Technology Projects of Gansu Province, and Lanzhou City’s scientific research funding subsidy to Lanzhou University.