### v2

# Charged Spherically Symmetric and Slowly Rotating Charged Black Hole Solutions in Bumblebee Gravity

Abstract
In this paper, we study the scenario in which the matter field is an electromagnetic field nonminimally coupled to the bumblebee vector field. We present exact charged spherically symmetric black hole solutions and slowly rotating charged solutions in bumblebee gravity with and without a cosmological constant. The static spherically symmetric solutions describe the Reissner-Nordström-like black hole and Reissner-Nordström-(anti) de Sitter-like black hole, while the stationary and axially symmetric soltuions describe the Kerr-Newman-like black hole and Kerr-Newman-(anti) de Sitter-like black hole. We utilize the Hamilton-Jacobi formalism to study the shadows of the black holes. Additionally, we investigate the effect of the electric charge and Lorentz-violating parameters on the radius of the shadow reference circle and the distortion parameter. We find that the radius of the reference circle decreases with the Lorentz-violating parameter and the charge parameter, while the distortion parameter increases with the Lorentz-violating parameter and the charge parameter.
Black Holes, Lorentz Symmetry Breaking,Black Hole Shadow
I INTRODUCTION
In the area of classical physics, gravitational phenomena are well-explained by general relativity, which has endured rigorous experimental and theoretical validation. On the quantum front, the Standard Model of particle physics has been successfully described the other three fundamental interactions. These two theories offer a comprehensive and understanding of the natural world. Therefore, the quest to unify these two theories is fundamental, and its achievement will inevitably lead us to a deeper comprehension of nature.
Several quantum gravity theories have been proposed Birrell and Davies (1984); Maldacena (1998); Aharony et al. (2000); Gubser et al. (1998); Alfaro et al. (2000, 2002); Rovelli and Smolin (1990), but it is needed the experiments at the Planck scale ($\sim 10^{19}$ GeV) to observe the quantum effect, this is beyond the ability of current experiments. However, some quantum gravity theories assume that Lorentz symmetry might be broken in the gravitational UV regime which is at our presently accessible low-energy scales. Stemming from the concept of spontaneous Lorentz symmetry breaking (LSB) in string theory Kostelecky and Samuel (1989a); Kostelecky (2004), the Standard Model extension was introduced as a suitable framework for violating Lorentz symmetry Colladay and Kostelecky (1997, 1998); Kostelecky (2004).In the following decades, the Standard Model extension has been widely studied Colladay and Kostelecky (2001); Kostelecky and Lehnert (2001); Kostelecky and Mewes (2001); Colladay and McDonald (2009); Berger et al. (2016); Carroll et al. (1990); Andrianov and Soldati (1995, 1998); Lehnert and Potting (2004); Baeta Scarpelli (2012); Brito et al. (2013).
In 1989, Kostelecky and Samuel introduced the prototype of the bumblebee model Kostelecky and Samuel (1989b, a), a string-inspired framework featuring tensor-induced spontaneous Lorentz symmetry breaking. In the framework of bumblebee gravity, the spontaneous Lorentz violation arises from a potential $V(B^{\mu}B_{\mu})$ acting on an vector field $B_{\mu}$ Kostelecky (2004); Kostelecky and Tasson (2011). Not just in Minkowski’s spacetime, the bumblebee model can also be explored in Riemann and Riemann-Cartan spacetimes Bailey and Kostelecky (2006); Maluf et al. (2013); Li and Övgün (2020); Bluhm et al. (2008a); Escobar and Martín-Ruiz (2017); Maluf et al. (2015); Hernaski (2014); Carroll et al. (2009).
The pursuit of black hole solutions holds significance for any theory of gravity. In 2017, Casana et al. obtained an exact Schwarzschild-like black hole solution in the bumblebee gravity Casana et al. (2018). Subsequently, researchers have independently proposed exact solutions for traversable wormholes Övgün et al. (2019), Schwarzschild-anti-de Sitter (AdS)-like black holes Maluf and Neves (2021), slowly rotating Kerr-like black holes Ding et al. (2020), and other solutions Santos et al. (2015); Jha and Rahaman (2021); Filho et al. (2023); Xu et al. (2023); Ding et al. (2023) within bumblebee gravity. Moreover, in other scenarios within the framework of the Standard Model Extension, such as when considering the nonminimally coupled Kalb-Ramond field with a nonzero vacuum expectation value, certain black hole solutions have also been obtained Lessa et al. (2020); Kumar et al. (2020); Yang et al. (2023); Duan et al. (2023); Do and Kao (2020); Liu et al. (2024a). And the properties of the blakc holes were also studied Guo et al. (2023); Du et al. (2024); Kuang and Övgün (2022); Oliveira et al. (2019); Liu et al. (2023); Gomes et al. (2020); Kanzi and Sakallı (2019); Güllü and Övgün (2022); Oliveira et al. (2021); Kanzi and Sakallı (2021); Hosseinifar et al. (2024); Liu et al. (2024b, a); Ovgün et al. (2018); Sakallı and Yörük (2023); Mangut et al. (2023); Uniyal et al. (2023).
In 1966, Synge initiated the examination of the shadow cast by a Schwarzschild black hole Synge (1966). Subsequently, Luminet Luminet (1979) studied the influence of a thin accretion disk on the black hole shadow, while Bardeen Bardeen (1973) investigated the shadow of a Kerr black hole. Over the decades, significant progress has been made in the study of black hole shadows de Vries (1999); Takahashi (2005); Bambi et al. (2012); Atamurotov et al. (2013); Perlick et al. (2018); Meng et al. (2023); Sui et al. (2023). Recently, direct imaging of the black holes M87* at the core of the Virgo A galaxy and Sgr A* at the core of the Milky Way galaxy has been achieved by the sub-millimeter Event Horizon Telescope (EHT) employing very-long baseline interferometry Akiyama et al. (2019a, b, c, d); Akiyama et al. (2022). Therefore, investigating the black hole shadow is highly desirable.
Our aim in this paper is to obtain Charged spherically symmetric and slowly rotating charged black hole solutions. Subsequently, using the equations governing the motion of photons, we analyze the shadows cast by rotating black holes, exploring the impact of Lorentz violation on the size and deformation of these shadows. Calculations are then conducted to determine the radius and distortion observables.
This paper is structured as follows: Sec. II provides a review of the Einstein-Bumblebee theory. In Sec. III, we present the charged bumblebee black hole solution in both the presence and absence of a cosmological constant. And in Sec. IV, we extend this solution to the case of rotating black holes. Sec. V is devoted to investigating the shadow of rotating black holes and computing relevant parameters. Finally, Sec. VI presents a summary and discussion of the work.
II EINSTEIN-BUMBLEBEE THEORY
As discussed in the preceding section, the bumblebee model extends general relativity by adding a vector field called bumblebee field which couples with gravity nonminimally. The bumblebee vector field $B_{\mu}$ obtains a nonzero vacuum expectation value through a specified potential, resulting in the spontaneous breaking of Lorentz symmetry in the gravitational sector. The action is described by  Kostelecky (2004)
$S$ $=$ $\int d^{4}x\sqrt{-g}\bigg{[}\frac{1}{2\kappa}\left(R-2\Lambda\right)+\frac{\xi}{2\kappa}B^{\mu}B^{\nu}R_{\mu\nu}-\frac{1}{4}B_{\mu\nu}B^{\mu\nu}-V(B^{\mu}B_{\mu}\pm b^{2})\bigg{]}$ (1)
$+\int d^{4}x\sqrt{-g}\mathcal{L}_{M},$
where $\Lambda$ denotes the cosmological constant, $\kappa=\frac{8\pi G}{c^{4}}$ is the gravitational coupling constant, and $\xi$ represents the nonminimal coupling constant between gravity and the bumblebee field. Analogous to the electromagnetic field, the bumblebee field strength is given by
$B_{\mu\nu}=\partial_{\mu}B_{\nu}-\partial_{\nu}B_{\mu}.$ (2)
The potential $V$ provides a nonvanishing vacuum expectation value for $B_{\mu}$ and could take the functional form of $V(B^{\mu}B_{\mu}\pm b^{2})$, implying that the condition $B^{\mu}B_{\mu}=\mp b^{2}$ must be satisfied. In essence, the field $B_{\mu}$ acquires a nonvanishing vacuum expectation value $\left\langle B_{\mu}\right\rangle=b_{\mu}$, where the vector $b_{\mu}$ is a function of the spacetime coordinates and satisfies $b^{\mu}b_{\mu}=\mp b^{2}=\mathrm{const}$. For convenience, we define $X=B^{\mu}B_{\mu}\pm b^{2}$ and $V^{\prime}=\frac{\partial V}{\partial X}$, which will be used in the subsequent discussions.
We consider the matter field to be an electromagnetic field nonminimally coupled with the bumblebee vector field Lehum et al. (2024). Its Lagrangian density expression reads as
$\mathcal{L}_{M}$ $=$ $\frac{1}{2\kappa}(F^{\mu\nu}F_{\mu\nu}+{\gamma}B^{\mu}B_{\mu}F^{\alpha\beta}F_{\alpha\beta}),$ (3)
where the electromagnetic tensor related to the electromagnetic field is
$F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu},$
$A=(\phi(r),0,0,0),$ (4)
and ${\gamma}$ presents the coupling coefficient. The gravitational field equations in the framework of bumblebee gravity can be derived by varying the action (1) with respect to the metric tensor $g^{\mu\nu}$:
$G_{\mu\nu}+\Lambda g_{\mu\nu}=\kappa T^{B}_{\mu\nu}+\kappa T^{M}_{\mu\nu},$ (5)
where
$T^{B}_{\mu\nu}$ $=$ $\frac{\xi}{\kappa}\left[\frac{1}{2}B^{\alpha}B^{\beta}R_{\alpha\beta}g_{\mu\nu}-B_{\mu}B^{\alpha}R_{\alpha\nu}-B_{\nu}B^{\alpha}R_{\alpha\mu}\right.+\frac{1}{2}\nabla_{\alpha}\nabla_{\mu}\left(B^{\alpha}B_{\nu}\right)$ (6)
$+\frac{1}{2}\nabla_{\alpha}\nabla_{\nu}\left(B^{\alpha}B_{\mu}\right)\left.-\frac{1}{2}\nabla^{2}\left(B_{\mu}B_{\nu}\right)-\frac{1}{2}g_{\mu\nu}\nabla_{\alpha}\nabla_{\beta}\left(B^{\alpha}B^{\beta}\right)\right]$
$+2V^{\prime}B_{\mu}B_{\nu}+B_{\mu}^{\ \alpha}B_{\nu\alpha}-\left(V+\frac{1}{4}B_{\alpha\beta}B^{\alpha\beta}\right)g_{\mu\nu},$
and
$T^{M}_{\mu\nu}=\frac{1}{\kappa}\left[(1+\gamma b^{2})(2F_{\mu\alpha}F_{\nu}^{\alpha}-\frac{1}{2}g_{\mu\nu}F^{\alpha\beta}F_{\alpha\beta})+\gamma B_{\mu}B_{\nu}F^{\alpha\beta}F_{\alpha\beta}\right].$ (7)
For computational convenience, we can express the field equations in the following form:
$R_{\mu\nu}=\Lambda g_{\mu\nu}+\kappa\mathcal{T}_{\mu\nu},$ (8)
where
$\mathcal{T}^{M}_{\mu\nu}$ $=$ ${T}^{M}_{\mu\nu}-\frac{1}{2}{T}^{M}g_{\mu\nu}$
$\mathcal{T}^{B}_{\mu\nu}$ $=$ ${T}^{B}_{\mu\nu}-\frac{1}{2}{T}^{B}g_{\mu\nu}$ (9)
and
$\kappa\mathcal{T}_{\mu\nu}$ $=$ $\kappa\mathcal{T}^{M}_{\mu\nu}+\kappa\mathcal{T}^{B}_{\mu\nu}$ (10)
$=$ $\kappa\left(T^{M}_{\mu\nu}-\frac{1}{2}g_{\mu\nu}T^{M}\right)+\kappa\left[V^{\prime}\left(2B_{\mu}B_{\nu}-b^{2}g_{\mu\nu}\right)+B_{\mu}^{\ \alpha}B_{\nu\alpha}+Vg_{\mu\nu}-\frac{1}{4}B_{\alpha\beta}B^{\alpha\beta}g_{\mu\nu}\right]$
$+\xi\left[\frac{1}{2}B^{\alpha}B^{\beta}R_{\alpha\beta}g_{\mu\nu}-B_{\mu}B^{\alpha}R_{\alpha\nu}-B_{\nu}B^{\alpha}R_{\alpha\mu}+\frac{1}{2}\nabla_{\alpha}\nabla_{\mu}\left(B^{\alpha}B_{\nu}\right)\right.+\frac{1}{2}\nabla_{\alpha}\nabla_{\nu}\left(B^{\alpha}B_{\mu}\right)$
$\left.-\frac{1}{2}\nabla^{2}\left(B_{\mu}B_{\nu}\right)\right].$
Similarly, by varying the action (1) with respect to the bumblebee vector field and the electromagnetic field, we can obtain the equations of motion for the corresponding fields:
$\nabla_{\mu}B^{\mu\nu}-2\left(V^{\prime}B^{\nu}-\frac{\xi}{2\kappa}B_{\mu}R^{\mu\nu}-\frac{1}{2\kappa}\gamma B^{\nu}F^{\alpha\beta}F_{\alpha\beta}\right)$ $=$ $0,$ (11)
$\nabla_{\mu}\left(F^{\mu\nu}+\gamma B^{\alpha}B_{\alpha}F^{\mu\nu}\right)$ $=$ $0$ (12)
III CHARGED SPHERICALLY SYMMETRIC BLACK HOLE SOLUTIONS
We consider the metric ansatz for a static and spherically symmetric spacetime, which is expressed as
${d}{s}^{2}=-A(r){dt}^{2}+S(r){dr}^{2}+r^{2}{\leavevmode\nobreak\ d}\Omega^{2},$ (13)
where ${\leavevmode\nobreak\ d}\Omega^{2}={\leavevmode\nobreak\ d}\theta^{2}+\sin^{2}\theta{d}\varphi^{2}$.
In the previous study conducted by Casana et al., they formulated an exact black hole solution without a cosmological constant and the matter field $A_{\mu}$ Casana et al. (2018). In the scenario where the bumblebee field $B_{\mu}$ maintains its vacuum expectation value $b_{\mu}$, the bumblebee field is given by Bertolami and Paramos (2005)
$B_{\mu}=b_{\mu}.$ (14)
Similar to Ref. Casana et al. (2018), we consider a spacelike background $b_{\mu}$ with the form
$b_{\mu}=\left(0,b_{r}(r),0,0\right).$ (15)
Utilizing the aforementioned condition $b_{\mu}b^{\mu}=b^{2}=const$, we can derive
$b_{r}(r)=b\sqrt{S(r)}.$ (16)
For convenience, we set $\ell=\xi b^{2}$ as Lorentz-violating parameter. By combining the spherically symmetric metric, we can obtain the specific field equations:
$\frac{(2+\ell)A^{\prime}(r)^{2}}{8A(r)S(r)}-\frac{(2+\ell)A^{\prime}(r)S^{\prime}(r)}{8S(r)^{2}}-\frac{(1+\ell)A^{\prime}(r)}{rS(r)}+\frac{(2+\ell)A^{\prime\prime}(r)}{4S(r)}$ (17)
$-\frac{\ell A(r)S^{\prime}(r)}{2rS(r)^{2}}-(1+2b^{2}\gamma)A(r)\phi^{\prime}(r)^{2}+\Lambda A(r)+\kappa(V^{\prime}b^{2}+V)A(r)$ $=$ $0,$ (18)
$\frac{A^{\prime}(r)S^{\prime}(r)}{4A(r)S(r)}-\frac{A^{\prime\prime}(r)}{2A(r)}+\frac{A^{\prime}(r)^{2}}{4A(r)^{2}}+\frac{S^{\prime}(r)}{rS(r)}$
$+\frac{2(1+2b^{2}\gamma)\phi^{\prime}(r)^{2}S(r)}{(2+3\ell)}-\frac{2\Lambda S(r)}{(2+3\ell)}-2\kappa\frac{(V^{\prime}b^{2}+V)S(r)}{(2+3\ell)}$ $=$ $0,$ (19)
$1-\frac{1+\ell}{S(r)}+\frac{rS^{\prime}(r)}{2S(r)^{2}}+\frac{\ell r^{2}A^{\prime\prime}(r)}{4A(r)S(r)}-\frac{\ell r^{2}A^{\prime}(r)S^{\prime}(r)}{8A(r)S(r)^{2}}-\frac{\ell r^{2}A^{\prime}(r)^{2}}{8A(r)^{2}S(r)}$
$-\frac{(1+\ell)rA^{\prime}(r)}{2A(r)S(r)}-(1+2b^{2}\gamma)\phi^{\prime}(r)^{2}r^{2}-\Lambda r^{2}-\kappa(V^{\prime}b^{2}+V)r^{2}$ $=$ $0,$ (20)
$\frac{A^{\prime\prime}(r)}{2A(r)}-\frac{A^{\prime}(r)^{2}}{4A(r)^{2}}-\frac{A^{\prime}(r)S^{\prime}(r)}{4A(r)S(r)}-\frac{2\gamma\phi^{\prime}(r)^{2}}{\xi\left(1+\ell\right)A(r)}-\frac{S^{\prime}(r)}{rS(r)}+\frac{2\kappa}{\xi}S(r)V^{\prime}$ $=$ $0,$ (21)
$\frac{d}{dr}\left[r^{2}\frac{F_{tr}}{\sqrt{A(r)S(r)}}\right]$ $=$ $0$ (22)
Next, we will solve the above equations and give spherically symmetric black hole solutions under different scenarios.
III.1 Case A: $V(X)=\frac{\lambda}{2}X^{2}$ and $\Lambda=0$
In the absence of the cosmological constant, similar to the work of Casana et al., we impose the vacuum conditions $V=0$ and $V^{\prime}=0$ Casana et al. (2018). An illustrative example of the potential satisfying these conditions is readily presented by a smooth quadratic form:
$V(X)=\frac{\lambda}{2}X^{2},$ (23)
where $\lambda$ is a constant.This form is same as the potential form of the Higgs field, and is also related to the mass structure of the theory Kostelecky and Samuel (1989b). In this case, the potential $V$ has no contribution to the field equations. Other choices of the potential, such as $V(X)\equiv\frac{\lambda}{2}X^{n}$ ($n\geq 3$), also have no contribution to the field equations. Therefore, the solutions for these choices are consistent with the potential $V(X)=\frac{\lambda}{2}X^{2}$.
Using Eq. (22), we can derive $F_{tr}=\sqrt{A(r)S(r)}\phi^{\prime}(r)$. Substituting the expressions of $F_{tr}$ here and $V$ in Eq. (23) into above conditions into the field equations (18) - (22), and considering the value of the coupling parameter $\gamma$ to be $\frac{\xi}{2+\ell}$, the field equations are given by:
$\frac{(2+\ell)A^{\prime}(r)^{2}}{8A(r)S(r)}-\frac{(2+\ell)A^{\prime}(r)S^{\prime}(r)}{8S(r)^{2}}-\frac{(1+\ell)A^{\prime}(r)}{rS(r)}$
$+\frac{(2+\ell)A^{\prime\prime}(r)}{4S(r)}-\frac{\ell A(r)S^{\prime}(r)}{2rS(r)^{2}}-\frac{A(r)\phi^{\prime}(r)^{2}}{\left(1+\ell\right)}$ $=$ $0,$ (24)
$\frac{A^{\prime}(r)S^{\prime}(r)}{4A(r)S(r)}-\frac{A^{\prime\prime}(r)}{2A(r)}+\frac{A^{\prime}(r)^{2}}{4A(r)^{2}}+\frac{S^{\prime}(r)}{rS(r)}+\frac{2\phi^{\prime}(r)^{2}S(r)}{(2+\ell)}$ $=$ $0,$ (25)
$1-\frac{1+\ell}{S(r)}+\frac{rS^{\prime}(r)}{2S(r)^{2}}+\frac{\ell r^{2}A^{\prime\prime}(r)}{4A(r)S(r)}-\frac{\ell r^{2}A^{\prime}(r)S^{\prime}(r)}{8A(r)S(r)^{2}}-\frac{\ell r^{2}A^{\prime}(r)^{2}}{8A(r)^{2}S(r)}$
$-\frac{(1+\ell)rA^{\prime}(r)}{2A(r)S(r)}-\frac{(2+3\ell)\phi^{\prime}(r)^{2}r^{2}}{(2+\ell)}$ $=$ $0,$ (26)
$\frac{A^{\prime\prime}(r)}{2A(r)}-\frac{A^{\prime}(r)^{2}}{4A(r)^{2}}-\frac{A^{\prime}(r)S^{\prime}(r)}{4A(r)S(r)}-\frac{\phi^{\prime}(r)^{2}}{\left(1+\ell\right)A(r)}-\frac{S^{\prime}(r)}{rS(r)}$ $=$ $0,$ (27)
$\frac{d}{dr}[r^{2}\phi^{\prime}(r)]$ $=$ $0$ (28)
By simplifying Eqs. (24) and (25), we obtain:
$[S(r)A(r)]^{\prime}=0,$ (29)
which implies that $S(r)=\frac{C_{1}}{A(r)}$, where $C_{1}$ is a constant. Similar to the Schwarzschild-like black hole solution in bumblebee gravity, we set the value of $C_{1}$ to be $1+\ell$ Casana et al. (2018). Building upon this, we solve the modified Maxwell’s equations to obtain: $\phi(r)=\frac{C_{2}}{r}+C_{3}$. Here, we set $C_{2}=Q_{0}$ and $C_{3}=0$ (i.e., $\phi(r)=\frac{Q_{0}}{r}$), and we ultimately arrive at the final solution:
$A(r)$ $=$ $1-\frac{2M}{r}+\frac{2(1+\ell)Q_{0}^{2}}{(2+\ell)r^{2}},$ (30)
$S(r)$ $=$ $\frac{1+\ell}{A(r)},$ (31)
$\phi(r)$ $=$ $\frac{Q_{0}}{r}.$ (32)
Through the modified Maxwell’s equations, the conserved current is modified to be $J^{\nu}=\nabla_{\mu}\left(F^{\mu\nu}+\gamma B^{\alpha}B_{\alpha}F^{\mu\nu}\right)$. Then, we can obtain the integration constants, the Maxwell’s equations was also modified.
[…]
$=$ $\left(1+b^{2}\frac{\xi}{(l+2)}\right)Q_{0}$
$=$ $\frac{2(1+\ell)}{2+\ell}Q_{0}.$
Here $\Sigma$ represents a three-dimensional spacelike region with the induced metric $\gamma_{ij}^{(3)}$, the boundary $\partial\Sigma$ is a two-sphere located at spatial infinity with the induced metric $\gamma_{ij}^{(2)}=r^{2}{\leavevmode\nobreak\ d}\Omega^{2}$, while $n_{\mu}=(1,0,0,0)$ and $\sigma_{\mu}=(0,1,0,0)$ denote the unit normal vectors associated with $\partial\Sigma$ and $\Sigma$, respectively. The solution of $A(r)$ can also be rewritten as
$A(r)=1-\frac{2M}{r}+\frac{(2+\ell)Q^{2}}{2(1+\ell)r^{2}},$ (34)
this solution is similar to the Reissner-Nordström (RN) solution. As $\ell$ tends to 0, it recovers to the RN solution. The introduction of Lorentz symmetry breaking distinguishes it from many modified gravity theories. It is evident that the metric functions approach to $A(r)\rightarrow 1$ and $S(r)\rightarrow 1+\ell$ at infinity, indicating that the spacetime is not asymptotically Minkowski.Furthermore, the $g_{tt}$ component of the metric (13) differs from the RN black hole solution only by the modification of the coefficient of the $\frac{Q^{2}}{r^{2}}$ term. However, the electrically charged black hole solutions in many modified gravity theories modify the terms involving both $\frac{M}{r}$ and $\frac{Q^{2}}{r^{2}}$ in the RN metric, or introduce new terms. For example, the charged black hole solutions in Einstein-Gauss-Bonnet gravity Fernandes (2020) and Eddington-inspired Born–Infeld gravity Wei et al. (2015a) modify the terms involving both $\frac{M}{r}$ and $\frac{Q^{2}}{r^{2}}$, while the charged black hole solution in Einstein-Maxwell-aether gravity introduces new term Ding et al. (2015). In the paper Casana et al. (2018), Casana et al. studied the effects of spontaneous Lorentz symmetry breaking and investigated some classic tests, including the advance of Mercury’s perihelion, bending of light, and Shapiro’s time-delay. Furthermore, they have computed several upper bounds, with the most stringent sensitivity of $\ell<6.2\times 10^{-13}$.
The introduction of the nonminimal coupling parameter $\ell$ does have some effects on black hole thermodynamics, such as the black hole temperature, entropy, and so on. However, it does not fundamentally alter the number of possible black hole phases, the phase transition behaviors, and the thermodynamic stability. In Appendix A, we have provided a series of specific analyses and discussions.
Similar to the RN black hole, this solution exhibits two horizons:
$r_{\pm}=M\pm\sqrt{M^{2}-\frac{2(\ell+1)}{\ell+2}Q_{0}^{2}}.$ (35)
It is obvious that the mass and charge parameter of the black hole should satisfy
$\frac{Q_{0}^{2}}{M^{2}}\leq\frac{(2+\ell)}{2(1+\ell)}.$ (36)
Another aspect of this solution is the physical singularity which can be studied through the Kretschmann scalar $K$, derived from the Riemann tensor:
$K$ $=$ $R_{\alpha\beta\delta\gamma}R^{\alpha\beta\delta\gamma}$ (37)
$=$ $\frac{4\left(\ell^{2}r^{2}+4lMr+12M^{2}\right)}{(\ell+1)^{2}r^{6}}-\frac{16Q_{0}^{2}\left((\ell+2)r(\ell r+12M)-14(\ell+1)Q_{0}^{2}\right)}{(\ell+1)(\ell+2)^{2}r^{8}}.$ (38)
From this we can see that, the singularity only appears at $r=0$ (assuming $0<\left|\ell\right|\ll 1$) which is similar to the singularity of the RN black hole.
To further clarify and describe the spacetime structure of the Charged black hole in Bumblebee gravity, we will present the Penrose diagram. By performing a coordinate transformation, we obtain:
${d}{s}^{2}=-(1-R_{+})(1-R_{-}){dt}^{2}+\frac{1}{(1-R_{+})(1-R_{-})}{dR}^{2}+\frac{1}{(1+\ell)}R^{2}{\leavevmode\nobreak\ d}\Omega^{2},$ (39)
Figure 1: Penrose diagram of the non-extremal charged black hole in bumblebee gravity.
where $R$ is the parameter after coordinate transformation of $r$ ($r=\sqrt{\frac{1}{1+\ell}}R$) and $R_{\pm}$ are the result of the two horizon radii after the coordinate transformation ($R_{\pm}=\sqrt{\frac{1}{1+\ell}}(M\pm\sqrt{M^{2}-\frac{2(\ell+1)}{\ell+2}Q_{0}^{2}})$). The corresponding Penrose diagrams are illustrated in Fig. 1. It is evident that the Penrose diagram of this solution is identical to that of the RN spacetime. The number and type of horizons of the electrically charged black holes obtained in our paper are indeed consistent with the RN black hole. The speed of light remains unchanged, and the causal structure is unaffected. However, due to Lorentz symmetry breaking in bumblebee gravity, the asymptotic behavior of the resulting spacetime is not Minkowskian. In addition, although the Penrose diagram obtained here is the same as that of the RN black hole, the two-dimensional spheres at the same points on the Penrose diagrams are not the same. Two horizon radii of the electrically charged black hole in bumblebee gravity differ from those of the RN black hole, and there is an additional correction factor 1($\frac{1}{(1+\ell)}$ in front of the $R^{2}d\Omega^{2}$ term compared to the RN black hole. It follows that, although our Penrose diagram is identical to that of the RN black hole after a series of coordinate transformations, the radii of the corresponding two-dimensional spheres in the diagrams are different.
III.2 Case B: $V(X)=\frac{\lambda}{2}X$ and $\Lambda\neq 0$
Next, we consider the scenario involving a nonzero cosmological constant. Our purpose is to derive an analytical black hole solution in this context. The geometry of the solution is similar to the RN-AdS black hole or the RN-dS black hole. Based on the groundwork by R.V. Maluf et al. Maluf and Neves (2021), conditions are different from those in the previous discussion. The straightforward option for the potential can be a linear function:
$V(X)=\frac{\lambda}{2}X,$ (40)
where $\lambda$ is a Lagrange multiplier Bluhm et al. (2008b). Note that the equation of motion for the Lagrange multiplier ensures the vacuum condition $X=0$, which also ensures that $V=0$ for any field $\lambda$ on-shell. However, unlike the vacuum condition in Case A, $V^{\prime}\neq 0$ is valid when the field $\lambda$ is nonzero. Thus, the potential $V$ contributes to the field equations. The choice $V(X)=\frac{\lambda}{2}X$ effectively freezes motion about the potential minimum and hence permits an efficient extraction of the essential physics Kostelecky and Samuel (1989b). The constraint for the Lagrange multiplier $\lambda$ is providedthe field Eqs (10) and (11) .
Combining the spherically symmetric metric and the form of the potential $V$, we can derive the specific field equations:
$\frac{(2+\ell)A^{\prime}(r)^{2}}{8A(r)S(r)}-\frac{(2+\ell)A^{\prime}(r)S^{\prime}(r)}{8S(r)^{2}}-\frac{(1+\ell)A^{\prime}(r)}{rS(r)}$
$+\frac{(2+\ell)A^{\prime\prime}(r)}{4S(r)}-\frac{\ell A(r)S^{\prime}(r)}{2rS(r)^{2}}-\frac{A(r)\phi^{\prime}(r)^{2}}{\left(1+\ell\right)}+\frac{\Lambda A(r)(2+\ell)}{2\ell}$ $=$ $0,$ (41)
$\frac{A^{\prime}(r)S^{\prime}(r)}{4A(r)S(r)}-\frac{A^{\prime\prime}(r)}{2A(r)}+\frac{A^{\prime}(r)^{2}}{4A(r)^{2}}+\frac{S^{\prime}(r)}{rS(r)}+\frac{2\phi^{\prime}(r)^{2}S(r)}{(2+\ell)}-\frac{\Lambda S(r)}{1+\ell}$ $=$ $0,$ (42)
$1-\frac{1+\ell}{S(r)}+\frac{rS^{\prime}(r)}{2S(r)^{2}}+\frac{\ell r^{2}A^{\prime\prime}(r)}{4A(r)S(r)}-\frac{\ell r^{2}A^{\prime}(r)S^{\prime}(r)}{8A(r)S(r)^{2}}-\frac{\ell r^{2}A^{\prime}(r)^{2}}{8A(r)^{2}S(r)}$
$-\frac{(1+\ell)rA^{\prime}(r)}{2A(r)S(r)}-\frac{(2+3\ell)\phi^{\prime}(r)^{2}r^{2}}{(2+\ell)}-\frac{\Lambda(2+\ell)r^{2}}{2(1+\ell)}$ $=$ $0,$ (43)
$\frac{A^{\prime\prime}(r)}{2A(r)}-\frac{A^{\prime}(r)^{2}}{4A(r)^{2}}-\frac{A^{\prime}(r)S^{\prime}(r)}{4A(r)S(r)}-\frac{\phi^{\prime}(r)^{2}}{\left(1+\ell\right)A(r)}-\frac{S^{\prime}(r)}{rS(r)}+\frac{\Lambda S(r)}{1+\ell}$ $=$ $0,$ (44)
$\frac{d}{dr}[r^{2}\phi^{\prime}(r)]$ $=$ $0$ (45)
An analytical solution is possible if and only if
$\Lambda=\frac{\kappa\lambda}{\xi}(1+\ell).$ (46)
Under the above condition, we can obtain
$A(r)$ $=$ $1-\frac{2M}{r}+\frac{2(1+\ell)Q_{0}^{2}}{(2+\ell)r^{2}}-\frac{1}{3}(1+\ell)\Lambda_{e}r^{2},$ (47)
$S(r)$ $=$ $\frac{1+\ell}{1-\frac{2M}{r}+\frac{2(1+\ell)Q_{0}^{2}}{(2+\ell)r^{2}}-\frac{1}{3}(1+\ell)\Lambda_{e}r^{2}},$ (48)
$\phi(r)$ $=$ $\frac{Q_{0}}{r}.$ (49)
For convenience, we set $\Lambda_{e}=\frac{\kappa\lambda}{\xi}$ as an effective cosmological constant, which satisfies the relation $R_{\mu\nu}=\Lambda_{e}g_{\mu\nu}$ at the boundary of the spacetime and is independent of the Lorentz-violating parameter.
Figure 2: Penrose diagram of the non-extremal charged AdS black hole in bumblebee gravity.
As $\ell$ approaches to 0, this solution becomes RN-(A)dS black hole solution. Similarly, we compute the Kretschmann scalar in order to study the singularity. The expression for $K$ is
$K$ $=$ $\frac{8}{3}\Lambda_{e}^{2}+\frac{l}{(l+1)r^{2}}\Lambda_{e}-\frac{16Q_{e}^{2}\left((\ell+2)r(\ell r+12M)-14(\ell+1)Q_{e}^{2}\right)}{(\ell+1)(\ell+2)^{2}r^{8}}$ (50)
$+$ $\frac{4\left(\ell^{2}r^{2}+4\ell Mr+12M^{2}\right)}{(\ell+1)^{2}r^{6}},$
which shows that, the RN-(A)dS-like singularity is only present at $r=0$ (assuming $0<\left|\ell\right|\ll 1$).
Through the coordinate transformation mentioned of $r$ ($r=\sqrt{\frac{1}{1+\ell}}R$), we obain
${d}{s}^{2}$ $=$ $-f(R){dt}^{2}+\frac{1}{f(R)}{dR}^{2}+\frac{1}{(1+\ell)}R^{2}{\leavevmode\nobreak\ d}\Omega^{2},$ (51)
where $f(R)=(1-\frac{2\sqrt{1+\ell}M}{R}+\frac{2(1+\ell)^{2}Q_{0}^{2}}{(2+\ell)R^{2}}-\frac{1}{3}\Lambda_{e}R^{2})$ then we can get the Penrose diagram of the non-extremal charged AdS black hole, as shown in Fig. 2, which is is identical to that of the classical RN-AdS spacetime. But the radii of the two-dimensional spheres at the same points on the Penrose diagrams are not the same.
IV SlOWLY ROTATING CHARGED BLACK HOLE SOLUTIONS
In this section, we will give slowly rotaing charged solutions through solving Einstein-bumblebee equations in both the cases with and without a cosmological constant. The forms of these solutions are very similar to the Kerr-Newman (KN) and Kerr-Newman (KN)-(A)dS solutions.
In 1963, Kerr discovered the rotating black hole solution Kerr (1963). A few years later, Newman obtained rotating charged black hole solution Newman et al. (1965). In bumblebee gravity, some rotating solutions were also provided Ding et al. (2020); Ding and Chen (2021); Jha and Rahaman (2021). Here, we consider slowly rotating charged black hole solution with and without a cosmological constant. The general form of the corresponding metric is given by
$ds^{2}=-A(r,\theta)dt^{2}+S(r,\theta)dr^{2}+2F(r)H(\theta)adtd\phi+\rho(r,\theta)^{2}d\theta^{2}+h(r,\theta)^{2}\sin^{2}\theta d\phi^{2}.$ (52)
The bumblee field and the $U(1)$ vector field are assumed as
$b_{\mu}$ $=$ $(0,b_{r}(r,\theta),0,0),$ (53)
$F_{\mu\nu}$ $=$ $\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu},$
$A_{\mu}$ $=$ $(A_{0}(r,\theta),0,0,A_{\phi}(r,\theta)),$ (54)
IV.1 Case A: $V(X)=\frac{\lambda}{2}X^{2}$ and $\Lambda=0$
It is noted that when the Lorentz-violating parameter is 0, this solution should reduce to the Kerr-Newman solution. Similarly, when $a=0$, we should obtain the spherically symmetric black hole solution derived in the previous section. Therefore, we consider the product of $a$ and $\ell$ is a small quantity.
Ding et al. obtained a slowly rotating black hole solution within bumblebee gravity Ding et al. (2020); Ding and Chen (2021). Following their approach and combining with Newman-Janis algorithm Newman and Janis (1965); Newman et al. (1965); Janis and Newman (1965), we obtained a slowly rotating charged black hole solution within bumblebee gravity. The metric in Boyer-Lindquist coordinates reads:
$ds^{2}$ $=$ $-\frac{\Delta_{r}}{\rho^{2}}\Big{(}\mathrm{d}t-a\sqrt{1+\ell}\sin^{2}\theta\mathrm{d}\phi\Big{)}^{2}+(1+\ell)\frac{\rho^{2}}{\Delta_{r}}\mathrm{d}r^{2}+{\rho^{2}}\mathrm{d}\theta^{2}$ (55)
$\quad+\frac{\sin^{2}\theta}{\rho^{2}}\Big{(}a\sqrt{1+\ell}\mathrm{d}t-(r^{2}+a^{2}(1+\ell))\mathrm{d}\phi\Big{)}^{2},$
$A_{0}(r,\theta)$ $=$ $-\frac{Q_{0}r}{\rho^{2}},$ (56)
$A_{\phi}(r,\theta)$ $=$ $\frac{Q_{0}a\sqrt{1+\ell}r\sin^{2}\theta}{\rho^{2}},$ (57)
$b_{r}(r)$ $=$ $b\rho\sqrt{\frac{\left(1+\ell\right)}{\Delta_{r}}},$ (58)
where
$\Delta_{r}$ $=$ $(1+\ell)a^{2}+r^{2}-2Mr+\frac{2(1+\ell)Q_{0}^{2}}{(2+\ell)},$
$\rho$ $=$ $\sqrt{r^{2}+(1+\ell)a^{2}\text{cos}^{2}\theta}.$ (59)
When $\ell\rightarrow 0$, it recovers to the usual Kerr-Newman metric and when $a\rightarrow 0$, it becomes to metric (13) with Eqs. (30) and (31). By solving $\Delta_{r}=0$, we can obatin the radii of the black hole’s horizon
$r_{\pm}=M\pm\sqrt{M^{2}-\frac{(1+\ell)(2Q_{0}^{2}+(2+\ell)a^{2})}{2+\ell}}.$ (60)
It becomes evident that the mass and spin parameters of the black hole must satisfy the following condition
$\frac{(1+\ell)(2Q_{0}^{2}+(2+\ell)a^{2})}{2+\ell}\leq M^{2}.$ (61)
Next, we discuss the small perturbations introduced by this rotating solution to the field equations, and impose certain constraints on the parameters through these perturbations. We set
$\Delta_{\mu\nu}=$ $R_{\mu\nu}$ $-\Lambda g_{\mu\nu}-\kappa\left[V^{\prime}\left(2B_{\mu}B_{\nu}+b^{2}g_{\mu\nu}\right)+B_{\mu}^{\ \alpha}B_{\nu\alpha}+Vg_{\mu\nu}\frac{1}{4}B_{\alpha\beta}B^{\alpha\beta}g_{\mu\nu}\right]$ (62)
-- $\kappa(T^{M}_{\mu\nu}-\frac{1}{2}g_{\mu\nu}T^{M})-\xi\left[\frac{1}{2}B^{\alpha}B^{\beta}R_{\alpha\beta}g_{\mu\nu}-B_{\mu}B^{\alpha}R_{\alpha\nu}-B_{\nu}B^{\alpha}R_{\alpha\mu}\right.$
$+$ $\frac{1}{2}\nabla_{\alpha}\nabla_{\mu}\left(B^{\alpha}B_{\nu}\right)+\frac{1}{2}\nabla_{\alpha}\nabla_{\nu}\left(B^{\alpha}B_{\mu}\right)\left.-\frac{1}{2}\nabla^{2}\left(B_{\mu}B_{\nu}\right)\right],$
naturally, for finite $Q_{0}$, when the field equations hold, we have $\Delta_{\mu\nu}=0$. Substituting the metric into the right hand side of Eq. (62), our calculations yield:
$\Delta_{\mu\nu}=\left\{\begin{aligned} \mathcal{O}(a^{2}\ell)&,&\quad\mu\nu=tt,rr,\phi\phi,\theta\theta,r\theta,t\phi\\ 0&,&\quad\text{otherwise}\end{aligned}\right.,$ (63)
At the same time, we derive the equations of motion for the bumblebee field and the electromagnetic field. Only the $r$ component nonvanishing:
$\left(\nabla_{\mu}B^{\mu r}-2(V^{\prime}B^{r}-\frac{\xi}{2\kappa}B_{\mu}R^{\mu r}-\gamma B^{r}F^{\alpha\beta}F_{\alpha\beta})\right)b$ $=$ $\mathcal{O}(a^{2}\ell),$
$\nabla_{\mu}\left(F^{\mu r}+\gamma B^{\alpha}B_{\alpha}F^{\mu r}\right)$ $=$ $\mathcal{O}(a^{2}\ell).$ (64)
Combining the above equations, we find that the rotating solution needs to satisfy the condition $\mathcal{O}(a^{2}\ell)$ being a small quantity. This implies that when considering this solution, $a$ and $\ell$ cannot both take large values simultaneously. Thus, for our solution, we need the condition that $a^{2}\ell$ is small quantities ($\frac{a^{2}\left|\ell\right|}{M^{2}}\ll 1$) for finite $Q_{0}$.
Theoretically, there is no constraint on the spin parameter $a$ and the Lorentz-violating parameter $\ell$. The constraint $a^{2}\left|\ell\right|\ll M^{2}$ or $a^{2}\left|\ell\right|\ll 1$ for $M=1$ on the parameters is imposed to ensure that our analytical solution approximately satisfies the field equations. We will use the analytical solution to study some observables of the black hole shadow, under the condition $a^{2}\left|\ell\right|\ll 1$. This constraint affects the range of the parameter combination $a^{2}\ell$, but it does not affect the individual parameter values. As $\ell$ becomes sufficiently small, $a$ may take larger values and even approach $M$. Conversely, when $a$ is sufficiently small, $\ell$ may take larger values. However, based on previous studies, it has been found that $\ell$ is constrained to a small value Casana et al. (2018), so in astrophysical scenarios, we consider $\ell$ to be sufficiently small, allowing a large range for $a$, which can even approach $M$.
When $\ell$ is sufficiently small, we can consider the extreme black hole case. The extreme black hole condition is approximately given by $\frac{(1+\ell)(2Q_{0}^{2}+(2+\ell)a^{2})}{2+\ell}=M^{2}$, since our solutions are not exact. For sufficiently small $\ell$, the extreme black hole condition can be approximated as $Q_{0}^{2}+a^{2}=M^{2}.$ From this, we can see that $a^{2}\leq M^{2}$. As long as $\ell$ is sufficiently small, $a^{2}\left|\ell\right|\ll M^{2}$ is still satisfied. In summary, provided that $\ell$ remains sufficiently small, our solution remains valid, even for the near-extremal KN-like black hole case.
IV.2 Case B: $V(X)=\frac{\lambda}{2}X$ and $\Lambda\neq 0$
Next, we consider the case of $V(X)=\frac{\lambda}{2}X$ and $\Lambda\neq 0$. Similar to the case A, we can get the following solution:
$ds^{2}$ $=$ $-\frac{\Delta_{r}}{\rho^{2}}\Big{(}\mathrm{d}t-\frac{a\sqrt{1+\ell}\sin^{2}\theta}{\Xi}\mathrm{d}\phi\Big{)}^{2}+(1+\ell)\frac{\rho^{2}}{\Delta_{r}}\mathrm{d}r^{2}+\frac{\rho^{2}}{\Delta_{\theta}}\mathrm{d}\theta^{2}$ (65)
$\quad+\frac{\sin^{2}\theta\Delta_{\theta}}{\rho^{2}}\Big{(}a\sqrt{1+\ell}\mathrm{d}t-\frac{r^{2}+a^{2}(1+\ell)}{\Xi}\mathrm{d}\phi\Big{)}^{2},$
$A_{0}(r,\theta)$ $=$ $-\frac{Q_{0}r}{\rho^{2}},$ (66)
$A_{\phi}(r,\theta)$ $=$ $\frac{Q_{0}ar\sin^{2}\theta}{\Xi\rho^{2}},$ (67)
$b_{r}(r)$ $=$ $b\rho\sqrt{\frac{\left(1+\ell\right)}{\Delta_{r}}},$ (68)
where
$\Delta_{\theta}$ $=$ $1+\frac{1}{3}(1+\ell)^{2}\Lambda_{e}a^{2}\text{cos}^{2}\theta,$
$\Delta_{r}$ $=$ $\left((1+\ell)a^{2}+r^{2}\right)\left(1-\frac{1}{3}(1+\ell)\Lambda_{e}r^{2}\right)-2Mr+\frac{2(1+\ell)Q_{0}}{(2+\ell)},$
$\Xi$ $=$ $1+\frac{1}{3}(1+\ell)^{2}\Lambda_{e}a^{2},$
$\rho$ $=$ $\sqrt{r^{2}+(1+\ell)a^{2}\text{cos}^{2}\theta}.$ (69)
From the form of the solution, it is a KN-(A)dS-like solution. Note that when the charge parameter $Q_{\text{0}}=0$, the solution is reduced to a slowly rotating Kerr-(A)dS-like black hole solution, which is not found in previous work. Otherwise, when $\ell=0$ and $a=0$, we get the usual KN-(A)dS black hole solution and the metric (13) with (47) and (48), respectively.
By calculating the errors in the field equations, we find that the results are similar to case A. Similarly, the rotating solution must satisfy the condition that $a^{2}\ell$ is a small quantity.
V SHADOW OBSERVABLES
The black hole shadow is formed by the strong gravitational field of the black hole. The shadow appears as a two-dimensional dark region to a distant observer. Shadows serve as one of the most intuitive pieces of evidence for the presence of black holes. Particularly, the unveiling of an image provides even stronger evidence for black hole’s existence Akiyama et al. (2019a, b, c, d); Akiyama et al. (2022). The shadow of the Kerr-like black hole in bumblebee gravity has been extensively studied in previous works Wang and Wei (2022); Ding et al. (2020); Kuang and Övgün (2022); Chen et al. (2020); Islam et al. (2024); Afrin et al. (2024). In this section, we aim to investigate the shadow of the KN-AdS like black holes in the framework of Einstein-bumblebee gravity.
V.1 The photon orbit
In this section, we will derive the equations of motion of photons orbiting the black holes. From Eqs. (63) and (64), it is apparent that the field equations hold when the rotating black hole solution satisfies that $a\ell^{2}$ is a small quantity. Therefore, in the subsequent part, all the analysies are under this condition. The motion of a photon in the background of the black hole is given by the geodesics. However, the process of solving geodesic equations is very difficult, we choose to use the Hamilton-Jacobi approach to solve it. The Lagrangian for the photon is
$\mathcal{L}=\frac{1}{2}g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu},$ (70)
where $\dot{x}$ is defined as $\dot{x}^{\mu}=dx^{\mu}/d\lambda=u^{\mu}$, with the four-velocity of the photon $u^{\mu}$ and the affine parameter $\lambda$. To facilitate the study of photon orbits and the black hole shadow observed by a distant observer, we introduce two parameters:
$\xi=\frac{L_{z}}{E}\,,\quad\eta=\frac{K}{E^{2}}\,,$ (71)
where $E$, $L_{z}$, and $K$ represent the energy, the axial component of the angular momentum, and the Carter constant Carter (1968), respectively. The expressions of the energy and the $z$-component of the angular momentum is given by
$E=-\frac{\partial{\mathcal{L}}}{\partial\dot{t}}=-g_{\varphi t}\dot{\varphi}-g_{tt}\dot{t},$ (72)
$L_{z}=\frac{\partial{\mathcal{L}}}{\partial\dot{\varphi}}=g_{\varphi\varphi}\dot{\varphi}+g_{\varphi t}\dot{t}.$ (73)
Using the Hamilton-Jacobi equation, we can get the null geodesic equation of the photon around the slowly rotating black hole:
$\dot{t}$ $=$ $\frac{E}{\rho^{2}}\Bigg{[}\frac{\left(r^{2}+(1+\ell)a^{2}\right)\left[\left(r^{2}+(1+\ell)a^{2}\right)-\sqrt{1+\ell}a\xi\Xi\right]}{\Delta_{r}}$
$+\frac{\sqrt{1+\ell}a\left(\xi\Xi-\sqrt{1+\ell}a\sin^{2}\theta\right)}{\Delta_{\theta}}\Bigg{]},$
$\dot{\varphi}$ $=$ $\frac{E\Xi}{\rho^{2}}\left[\frac{\sqrt{1+\ell}a(\left(r^{2}+(1+\ell)a^{2}\right)-\sqrt{1+\ell}a\xi\Xi)}{\Delta_{r}}+\frac{\xi\Xi-\sqrt{1+\ell}a\sin^{2}\theta}{\sin^{2}\theta\Delta_{\theta}}\right],$ (75)
$(1+\ell)(\rho^{2}\dot{r})^{2}$ $=$ $E^{2}\left[\left[\left(r^{2}+(1+\ell)a^{2}\right)-\sqrt{1+\ell}a\xi\Xi\right]^{2}-\Delta_{r}\Big{[}\frac{\left(\sqrt{1+\ell}a-\xi\Xi\right)^{2}}{\Delta_{\theta}}+\eta\Big{]}\right]$ (76)
$\equiv$ $(1+\ell)\mathcal{R}(r),$
$\rho^{2}\dot{\theta}$ $=$ $E^{2}\left[\eta\Delta_{\theta}-\cos^{2}\theta\left(\frac{\xi^{2}\Xi^{2}}{\sin^{2}\theta}-(1+\ell)a^{2}\right)\right]\equiv\Theta(\theta).$ (77)
Figure 3: The effective potential $V_{\text{eff}}$ for the photon. The paramengts are set as $M=1,a=0.1$ , $Q_{0}=0.1$ , $l=0,0.1$ and $\xi_{c}$ is an the value of $\xi$ for unstable circular null orbit.
Photons with sufficiently large orbital angular momentum can escape from the black hole, while those with small angular momentum will fall into the black hole. Now we focus on the critical case, the circular null orbit, where the photons will neither escape from nor fall into the black hole. The radial motion can also be expressed as
$(\rho^{2}\dot{r})^{2}+V_{\text{eff}}=0,$ (78)
where the effective potential $V_{\text{eff}}$ reads
$V_{\text{eff}}=-\frac{E^{2}}{(1+\ell)}\left[\left[\left(r^{2}+(1+\ell)a^{2}\right)-\sqrt{1+\ell}a\xi\Xi\right]^{2}-\Delta_{r}\big{[}\frac{\left(\sqrt{1+\ell}a-\xi\Xi\right)^{2}}{\Delta_{\theta}}+\eta\big{]}\right].$ (79)
In Fig. 3, we depict the effective potential $V_{\text{eff}}$ for a rotating charged black hole with $\ell=0$ and $\ell=0.1$. From Fig. 3, we can observe that Lorentz-violating parameter greater than 0 can increase the effective potential Wei et al. (2015b). For the unstable circular null orbit, the effective potential satisfies
${V}(r)=\frac{d{V}(r)}{dr}=0,\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \frac{d^{2}V(r)}{dr^{2}}<0,$ (80)
then we can obtain Tsukamoto et al. (2014); Tsukamoto (2018)
$\eta$ $=$ $\frac{r^{2}\left(16(1+\ell)a^{2}\Delta_{r}\Delta_{\theta}-\left(r\Delta_{r}^{\prime}-4\Delta_{r}\right)^{2}\right)}{(1+\ell)a^{2}\Delta_{\theta}\Delta_{r}^{\prime 2}},$ (81)
$\xi$ $=$ $\frac{\left(r^{2}+(1+\ell)a^{2}\right)\Delta_{r}^{\prime}-4r\Delta_{r}}{\sqrt{1+\ell}a\Xi\Delta_{r}^{\prime}}.$ (82)
(a)  $M=1,a=0.1,\ell=0.01,Q_{0}=0.2,\theta=\frac{\pi}{2}.$
(b)  $M=1,a=0.9,\ell=0.01,\Lambda_{e}=-0.1,Q_{0}=0.3.$
(c)  $M=1,a=0.1,\Lambda_{e}=-0.1,Q_{0}=0.2.$
(d)  $M=1,a=0.647,\ell=0.01,\Lambda_{e}=-0.2,\theta=\frac{\pi}{2}.$
Figure 4: The shadows for the KN-(A)dS-like black hole with different parameters.
V.2 Observables
We now aim to determine the apparent shape of the slowly rotating charged black hole’s shadow. The apparent shape of the black hole shadow for an observer far away from the black hole can be described using the celestial coordinates $x$ and $y$:
$x$ $=$ $\lim_{r\to\infty}\frac{-rp^{(\varphi)}}{p^{(t)}}=-\xi\csc\theta,$ (83)
$y$ $=$ $\lim_{r\to\infty}\frac{rp^{(\theta)}}{p^{(t)}}=\pm\sqrt{\eta+a^{2}\cos^{2}\theta-\xi^{2}\cot^{2}\theta}.$ (84)
The coordinate $x$ denotes the apparent perpendicular distance from the shape’s projection on the equatorial plane, while $y$ represents the apparent perpendicular distance of the shape from the axis of symmetry. The boundary curve of the shadow is determined by the unstable circular null orbit. Substituting the metric into Eqs. (81) and (82), we have
$\eta$ $=$ $\frac{4r^{2}\left(\left(a^{2}(\ell+1)+r^{2}\right)\left(1-\frac{1}{3}(\ell+1)r^{2}\Lambda_{e}\right)-2Mr+\frac{(1+\ell)Q_{0}^{2}}{(1+\frac{\ell}{2})}\right)}{\left(\frac{1}{3}(\ell+1)r\Lambda_{e}\left(a^{2}(\ell+1)+2r^{2}\right)+M-r\right)^{2}}$ (85)
$-\frac{r^{2}\left(-\frac{1}{3}a^{2}(\ell+1)r^{2}\Lambda_{e}+2a^{2}-\frac{3Mr}{1+\ell}+\frac{4Q_{0}^{2}}{(2+\ell)}+r^{2}\right)^{2}}{a^{2}\left(\frac{1}{3}a^{2}(\ell+1)^{2}\Lambda_{e}\cos^{2}(\theta)+1\right)\left(\frac{1}{3}(\ell+1)r\Lambda_{e}\left(a^{2}(\ell+1)+2r^{2}\right)+M-r\right){}^{2}},$
$\xi$ $=$ $\frac{\frac{1}{3}a^{2}(\ell+1)^{2}r\Lambda_{e}\left(a^{2}(\ell+1)+r^{2}\right)+a^{2}(\ell+1)(M+r)+r\left(r(r-3M)+\frac{4Q_{0}^{2}}{(2+\ell)}\right)}{a\sqrt{\ell+1}\left(\frac{1}{3}a^{2}(\ell+1)^{2}\Lambda_{e}+1\right)\left(\frac{1}{3}(\ell+1)r\Lambda_{e}\left(a^{2}(\ell+1)+2r^{2}\right)+M-r\right)}.$ (86)
(a)  $\theta=\frac{\pi}{2}$
(b)  $\theta=\frac{\pi}{4}$
(c)  $\theta=\frac{\pi}{6}$
Figure 5: The radius $R_{s}$ of the shadow against the Lorentz-violating parameter $\ell$.
(a)  $\theta=\frac{\pi}{2}$
(b)  $\theta=\frac{\pi}{4}$
(c)  $\theta=\frac{\pi}{6}$
Figure 6: The distortion parameter $\delta_{s}$ against the Lorentz-violating parameter $\ell$.
It is easy to see that both the coordinates $x$ and $y$ are the functions of $(M,Q_{eff},a,\Lambda_{e},\ell,\theta)$. We set $M=1$, and show the shapes of the shadow in Fig. 4. Similar to the KN-AdS black hole, the shadow radius increases with the negative effective cosmological costant $\Lambda_{e}$, as shown in Fig. 4(a). From Fig. 4(b) we can see that, the position of the shadow depends on the observation angle $\theta$. The Lorentz-violating parameter $\ell$ also affects the shaow size which can be seen from Fig. 4(c). Besides, the deformation of the shadow becomes more pronounced when the charge parameter $Q_{0}$ increases, which can be seen from Fig. 4(d).
(a)   $\theta=\frac{\pi}{2}$
(b)   $\theta=\frac{\pi}{4}$
(c)  $\theta=\frac{\pi}{6}$
Figure 7: The radius $R_{s}$ of the shadow against the charge parameter $Q_{0}$.
(a)  $\theta=\frac{\pi}{2}$
(b)  $\theta=\frac{\pi}{4}$
(c)  $\theta=\frac{\pi}{6}$
Figure 8: The distortion parameter $\delta_{s}$ against the charge parameter $Q_{0}$.
To effectively analyze astronomical observational data, it is essential to establish easily measurable and reliable astronomical observables. In this study, we will examine the following observables, the radius $R_{s}$ and the distortion parameter $\delta_{s}$ which are defined by Hioki and Maeda Hioki and Maeda (2009). The parameter $R_{s}$ represents the radius of the reference circle, which is expected to pass through the top point $(x_{t},y_{t})$, the bottom point $(x_{b},y_{b})$, and the right point $(x_{r},0)$ of the shadow. Meanwhile, the distortion parameter quantifies the distortion of the black hole shadow in comparison with the reference circle, and they can be expressed as
$R_{s}=\frac{(x_{t}-x_{r})^{2}+y_{t}^{2}}{2(x_{r}-x_{t})},\quad\delta_{s}=\frac{x_{l}-\tilde{x}_{l}}{R_{s}},$ (87)
where $(x_{l},0)$ and $(\tilde{x}_{l},0)$ represent the left points of the shadow and the reference circle, respectively.
For a given spin parameter, we illustrate the radius $R_{s}$ and the distortion parameter $\delta_{s}$ as functions of the Lorentz-violating parameter $\ell$ for $\theta=\frac{\pi}{2},\theta=\frac{\pi}{4},\theta=\frac{\pi}{6}$. The behavior of $R_{s}$ and $\delta_{s}$ is exhibited in Fig. 6 and Fig. 6. For fixed $a$ and $\theta$, the radius $R_{s}$ decreases with $\ell$, but $\delta_{s}$ increases with $\ell$ increases. Particularly noteworthy is the observation that for observers situated far from the equatorial plane, the shadow of a rapidly rotating black hole appears smaller compared to that of a slower one with the same $\ell$. However, this discrepancy diminishes as the observation angle increases, eventually resulting in a convergence of all curves as $\theta$ approaches $\frac{\pi}{2}$. Furthermore, as the observer progressively approaches the equatorial plane, the distortion of the shadow becomes increasingly pronounced.
Similarly, we investigate the relationship between the value of the charge parameter and $R_{s}$ and $\delta_{s}$ which is shown in Fig. 8 and Fig. 8. We find that $R_{s}$ decreases with the charge parameter, whereas $\delta_{s}$ increases with the charge parameter. Another notable observation in our study is that for small values of the spin parameter $a$, the deformation of the shadow is nearly imperceptible.
The shadow image of the supermassive black hole M87* was photographed by the EHT collaboration with the mass $M=6.5\times 10^{9}M_{\odot}$, the distance $r_{0}=16.8Mpc$, and the inclination angle $\theta_{0}=17^{o}$ Akiyama et al. (2019a, b, c, d); Akiyama et al. (2022). And in Ref. Kocherlakota et al. (2021), the authors stated that the shadow size of M87* should lie between $4.31M$ and $6.08M$. We will use the shadow observable $R_{a}$ to constrain the parameters of the Kerr-Newman-like black hole in bumblebee gravity. The definition of the characteristic areal radius of the shadow curve can be expressed as Banerjee et al. (2020); Abdujabbarov et al. (2015)
$R_{a}=\sqrt{\frac{2}{\pi}\int^{r_{max}}_{r_{min}}\Big{(}y(r)\frac{dx(r)}{dr}\Big{)}}.$ (88)
Figure 9: The constraints of the EHT observation $4.31M\leq R_{a}\leq 6.08M$ on the parameters ($a,Q_{0}$). The real curves represent $R_{a}=4.31M$ with different Lorentz-violating parameters $\ell$.
Figure 9 shows the constraints on the black hole parameters by the EHT observation with the shadow size $R_{a}$. All the solid lines represent the lower bound $R_{a}=4.31M$ or $\frac{(1+\ell)(2Q_{0}^{2}+(2+\ell)a^{2})}{2+\ell}=M^{2}$. The reasonable parameters, subject to the constraints, lie below these lines. The results illustrate that the larger the Lorentz-violating parameter $\ell$, the stronger the constraint on the rotating parameter $a$ and charge parameter $Q_{0}$. For $\ell=-0.1$ and $\ell=0.1$, the ranges of the charge paramter $Q_{0}$ are $Q_{0}/M\in(0,0.93)$ and $Q_{0}/M\in(0,0.88)$, respectively. When we consider a small Lorentz-violating parameter, according to the range of the shadow size $R_{a}$, we can not make any constraint on the rotating parameter $a$.
VI CONCLUSIONS
In this paper, by considering the coupling between the electromagnetic field and the bumblebee vector field, we obtained the RN-like black hole solutions and the RN-(A)dS-like black hole solutions in bumblebee gravity. When the Lorentz-violating parameter vanishes $(\ell=0)$, these solutions reduce to the RN and RN-(A)dS black hole solutions. In the stationary and axially symmetric scenario, we obtained solutions for slowly rotating charged black holes, i.e., the KN-(A)dS-like black hole, within the framework of bumblebee gravity. We further analyzed the series expansion of small quantities.
Because the nonminimal interaction term between the electromagnetic field and bumblebee field was introduced, the Maxwell’s equations was also modified. The charge term of the black hole solution is related to the Lorentz-violating parameter. When the electromagnetic field does not exist, our solutions are then reduced to another bumblebee black hole solutions in previous works.
Moreover, in the background of the KN-(A)dS-like black hole, we derived null geodesics using the Hamilton-Jacobi equation. Utilizing the conditions for unstable circular null orbits, we derived two conserved parameters, $\xi$ and $\eta$, which govern the photon’s motion. Using these parameters, we determined the celestial coordinates of the shadow. Subsequently, we investigated the size and deformation of the shadow. By varying the observation angle $\theta=\frac{\pi}{2},$ $\theta=\frac{\pi}{4},$ $\theta=\frac{\pi}{6}$, we calculated the shadow’s radius $R_{s}$ for different parameters. We observed that for fixed spin parameter $a$ and the Lorentz-violating parameter $\ell$, observers situated at lower latitudes consistently observe a larger shadow. Meanwhile, we found that the radius of the reference circle decreases with the Lorentz-violating parameter and electric charge , while the distortion parameter increases with the Lorentz-violating parameter and electric charge.
Acknowledgements
We are grateful to Tao-Tao Sui and Ke Yang for useful discussions. This work was supported by the National Key Research and Development Program of China (Grant No. 2021YFC2203003), the National Natural Science Foundation of China (Grants No. 12475056, No. 12247101, No. 12475055 and No. 12205129), the 111 Project (Grant No. B20063), Gansu Province’s Top Leading Talent Support Plan.
Appendix A A BRIEF DISCUSSION ON BLACK HOLE THERMODYNAMICS
In this appendix, we briefly discuss the impact of $\ell$ $(\ell=\xi b^{2})$ on black hole thermodynamics. First, using the surface gravity, the black hole temperature is given by
$T_{h}=\frac{-2{Q_{0}}^{2}(1+\ell)-{r_{h}}^{2}(2+\ell)\left(\Lambda_{e}{r_{h}}^{2}(1+\ell)-1\right)}{4\pi{r_{h}}^{3}\sqrt{1+\ell}(2+\ell)},$ (89)
where $r_{h}$ is the event horizon radius. Noting that when $\left|\ell\right|<1$, we can deduce that by redefining certain physical quantities ($Q=\sqrt{\frac{2(1+\ell)}{2+\ell}}{Q_{0}}$, $\Lambda=(1+\ell)\Lambda_{e}$, $T_{h}^{\prime}=\sqrt{1+\ell}T_{h}$), the black hole temperature can be restored to the original form of the RN-AdS black hole. This means that the curve of the temperature as a function of the horizon radius does not change fundamentally. For the RN-AdS-like black hole in bumblebee gravity, the extended first law of thermodynamics can be formulated as
$dM=T_{h}dS+\Phi_{0}dQ_{0}+\frac{\Theta}{8\pi}d\Lambda_{e},$ (90)
where $S=4\pi\sqrt{1+\ell}r_{h}^{2}$, $\Theta=-\frac{4}{3}(1+\ell)\pi r_{h}^{3}$, $\Phi_{0}=\frac{2(1+\ell)Q_{0}}{(2+\ell)r}$. This is similar to the first law of thermodynamics for the RN-AdS black hole.
Furthermore, by using the critical point equations
$\partial_{r_{h}}T\left(r_{h}\right)=0,\quad\partial_{r_{h}}^{2}T\left(r_{h}\right)=0,$ (91)
we can easily obtain $Q_{0}=\frac{\sqrt{2+\ell}r_{h}}{\sqrt{6(1+\ell)}}$ and $\Lambda_{e}=\frac{1}{2{r_{h}}^{2}(1+\ell)}$ which only differ from the critical values in the RN-AdS black hole by a coefficient. Therefore, the critical point has indeed been modified, but there is no essential difference from the RN-AdS black hole.
The specific heat can be calculated as
$C_{\text{B}}=T_{h}\left(\frac{\partial S}{\partial T_{h}}\right)_{Q_{0},\Lambda_{e}}=\frac{2\pi\sqrt{1+\ell}r_{h}^{2}\left(-\left(\Lambda_{e}(1+\ell)r_{h}^{4}\right)+r_{h}^{2}-\frac{2{Q_{0}}^{2}(1+\ell)}{\ell+2}\right)}{\frac{6{Q_{0}}^{2}(1+\ell)}{2+\ell}-r_{h}^{2}\left(\Lambda_{e}(1+\ell)r_{h}^{2}+1\right)}.$ (92)
It can be seen that the specific heat has been modified. By redefining certain physical quantities ($Q=\sqrt{\frac{2(1+\ell)}{2+\ell}}{Q_{0}}$, $\Lambda=(1+\ell)\Lambda_{e}$), the specific heat becomes equal to that of the RN-AdS black hole, multiplied by a factor, satisfying $C_{\text{B}}=\sqrt{1+\ell}C_{\text{R}}$. It is found that for $01+12Y>$ 0, where $Y=\frac{(1+\ell)^{2}{Q_{0}}^{2}\Lambda_{e}}{(1+\frac{\ell}{2})}$, the heat capacity is positive within the intervals $r_{h}\subset\left(\sqrt{\frac{\sqrt{1-4Y}-1}{-2(1+\ell)\Lambda_{e}}},\sqrt{\frac{1-\sqrt{1+12Y}}{-2(1+\ell)\Lambda_{e}}}\right)\cup\left(\sqrt{\frac{1+\sqrt{1+12Y}}{-2(1+\ell)\Lambda_{e}}},\infty\right)$. While $1+12Y\leq 0$, the heat capacity is positive for $r_{h}\subset\left(\sqrt{\frac{\sqrt{1-4Y}-1}{-2(1+\ell)\Lambda_{e}}},\infty\right)$. From this, we conclude that stability of black hole thermodynamics does not change fundamentally. On the other hand, from the black hole free energy
$F=M-T_{h}S=\frac{1}{4}\left(\frac{3(1+\ell)Q_{0}^{2}}{\left(1+\frac{\ell}{2}\right)r_{h}}+3(1+\ell)\Lambda_{e}r_{h}^{3}+r_{h}\right),$ (93)
it reveals the possible black hole phase transitions are similar to those of the RN-AdS black hole, the system may exhibit black hole phases with the following stability: small black holes (stable), medium black holes (unstable), and large black holes (stable).
From these aspects, we can conclude that the nonminimal coupling does indeed modify the black hole thermodynamics, but it does not change some of the essential behaviors. Additionally, the modification parameter is generally small, indicating that Lorentz symmetry breaking has a relatively minor impact on black hole thermodynamics in this model.